### v1

# Around the Merino–Welsh conjecture: improving Jackson’s inequality

Abstract.
The Merino-Welsh conjecture states that for a graph $G$ without loops and bridges we have
$\max(T_{G}(2,0),T_{G}(0,2))\geq T_{G}(1,1).$
Later Jackson proved that for any matroid $M$ without loop and coloop we have
$T_{M}(3,0)T_{M}(0,3)\geq T_{M}(1,1)^{2}.$
The value $3$ in this statement was improved to $2.9242$ by Beke, Csáji, Csikvári and Pituk. In this paper, we further improve on this result by showing that
$T_{M}(2.355,0)T_{M}(0,2.355)\geq T_{M}(1,1)^{2}.$
We also prove that the Merino–Welsh conjecture is true for matroids $M$, where all circuits of $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$ for some $\ell\geq 6$.
The research was supported by the MTA-Rényi Counting in Sparse Graphs ”Momentum” Research Group, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
1. Introduction
Let $G$ be a connected graph without loops and bridges. Merino and Welsh [18] conjectured that
$\max\left(\alpha(G),\alpha^{*}(G)\right)\geq\tau(G),$
where $\alpha(G),\alpha^{*}(G),\tau(G)$ denote the number of acyclic orientations, strongly connected orientations, and spanning trees of $G$, respectively. These quantities are evaluations of the Tutte polynomial $T_{G}(x,y)$, namely $T_{G}(2,0)=\alpha(G)$, $T_{G}(0,2)=\alpha^{*}(G)$ and $T_{G}(1,1)=\tau(G)$.
Conde and Merino [5] proposed "additive" and "multiplicative" versions of this conjecture:
$T_{G}(2,0)+T_{G}(0,2)$ $\geq 2T_{G}(1,1)$
$T_{G}(2,0)T_{G}(0,2)$ $\geq T_{G}(1,1)^{2},$
respectively. The multiplicative version implies the additive version, which in turn implies the original conjecture. These conjectures also naturally extend to the Tutte polynomial of a matroid without loops and coloops.
While the conjecture holds for certain classes of graphs [16, 19, 21] and matroids [4, 10, 14, 15, 17] it fails for general matroids. Beke, Csáji, Csikvári and Pituk [2] showed that there exist infinitely many matroids without loops and coloops violating the multiplicative version.
Theorem 1.1 (Beke, Csáji, Csikvári and Pituk [2]).
There are infinitely many matroids $M$ without loops and coloops for which
$T_{M}(2,0)T_{M}(0,2)<T_{M}(1,1)^{2}.$
Furthermore, let $x_{0}$ be the largest root of the polynomial $x^{3}-9(x-1)$. ($x_{0}\approx 2.22668...$) Then for $0\leqslant a<x_{0}$ there are infinitely many matroids $M$ without loops and coloops for which
$T_{M}(a,0)T_{M}(0,a)<T_{M}(1,1)^{2}.$
This paper investigates the following question: for what values of a does the inequality
$T_{M}(a,0)T_{M}(0,a)\geq T_{M}(1,1)^{2}$
hold for all loopless and coloopless matroids $M$? In this direction the first major result is due to Jackson.
Theorem 1.2 (Jackson [13]).
For any matroid M without loops and coloops,
$T_{M}(3,0)T_{M}(0,3)\geq T_{M}(1,1)^{2}.$
Jackson’s result was improved by Beke, Csáji, Csikvári and Pituk in the paper [1]. They showed that one can write $2.9243$ instead of $3$. In this paper, we further improve on this inequality.
Theorem 1.3.
For any matroid M without loops and coloops and $a\geq 2.355$,
$T_{M}(a,0)T_{M}(0,a)\geq T_{M}(1,1)^{2}.$
We may also study which matroid classes satisfy the product version of the Merino–Welsh conjecture. The following theorem is motivated by the fact that paving matroids satisfy the Merino–Welsh conjecture. A matroid of rank $r$ is a paving matroid if all circuits have length $r$ or $r+1$.
Theorem 1.4.
Suppose that there exists an $\ell\geq 6$ such that all circuits of the matroid $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$. Then
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
The proof of Theorem 1.3 and Theorem 1.4 are based on the theory of the permutation Tutte polynomial developed in the paper [1]. While Theorem 1.4 does not imply that paving matroids satisfy the Merino–Welsh conjecture, one can prove this fact by modifying the proof of Theorem 1.4.
Notations. Throughout the paper $G=(V,E)$ is an arbitrary graph and $H=(A,B,E)$ is a bipartite graph. $K_{a,b}$ denotes the complete bipartite graph with parts of size $a$ and $b$. $S_{k}$ denotes the star graph on $k$ vertices. For a vertex $v$ the degree of $v$ is denoted by $d_{v}$. $N_{H}(v)$ denotes the set of neighbors of $v$. If $H$ is clear from the context, then we simply write $N(v)$.
This paper is organized as follows.
  • In the next section we introduce the basic concepts from matroid theory that we will use, and revisit the theory of permutation Tutte polynomial $\widetilde{T}_{H}(x,y)$ developed in the paper [1].
  • In Section 3 we compute the growth constant $\lim_{n\to\infty}\widetilde{T}_{H_{n}}(x,0)^{1/n}$ for some family of bipartite graphs. While this section is not necessary for the proofs of Theorem 1.3 and 1.4 it provides an important intuition to study the permutation Tutte polynomial.
  • In Section 4 we prove a technical, but very important lemma that provides the basis of the proof of Theorems 1.3 and 1.4.
  • In Section 5 we give the proof of Theorem 1.3.
  • In Section 6 we prove Theorem 1.4.
  • In Section 7 we end the paper some conjectures.
  • In the Appendix one can find some tables that are used in the proof of Theorem 1.3.
2. Preliminaries
This section collects the necessary tools from matroid theory together, and recall some of the basic facts from the theory of permutation Tutte polynomial.
2.1. Tutte polynomial and matroids
The Tutte polynomial $T_{G}(x,y)$ of a graph $G$ is defined as
$T_{G}(x,y)=\sum_{A\subseteq E}(x-1)^{k(A)-k(E)}(y-1)^{k(A)+|A|-v(G)},$
with $k(A)$ denoting the number of connected components of the graph $(V,A)$, see [22]. There is a vast amount of literature on the properties of the Tutte polynomial and its applications, see for instance, [3, 6, 8, 23], or the book [9].
The Tutte polynomial naturally extends to matroids. Recall that a matroid $M$ is a pair $(E,\mathcal{I})$ such that $\mathcal{I}\subseteq 2^{E}$, called independent sets, satisfies the axioms (i) $\emptyset\in\mathcal{I}$, (ii) if $A^{\prime}\subseteq A\in\mathcal{I}$, then $A^{\prime}\in\mathcal{I}$, and (iii) if $A,B\in\mathcal{I}$ such that $|B|<|A|$, then there exists an $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{I}$. Given a set $S\subseteq E$, the maximal independent subsets of $S$ all have the same cardinality, and this cardinality is called the rank of the matroid, denoted by $r(S)$. The maximum size independent sets of $M$ are called bases, and their set is denoted by $\mathcal{B}(M)$. The dual of a matroid $M$ is the matroid $M^{*}$ whose bases are $\{E\setminus B\ |\ B\in\mathcal{B}(M)\}$. For further details on matroids, see for instance [20]
Given a graph $G=(V,E)$, the edge sets of the spanning forests of $G$ form the independent sets of a matroid $M_{G}$ called the cycle matroid of $G$. If $G$ is connected, then the basis of $M_{G}$ are the spanning trees of $G$. One can define the Tutte polynomial of a matroid as
$T_{M}(x,y)=\sum_{S\subseteq E}(x-1)^{r(E)-r(S)}(y-1)^{|S|-r(S)},$
where $r(S)$ is the rank of a set $S\subseteq E$. When $M=M_{G}$, then $T_{M_{G}}(x,y)=T_{G}(x,y)$. A loop in a matroid $M$ is an element $x\in E$ such that $r(\{x\})=0$, that is, $\{x\}\notin\mathcal{I}$, and a coloop is an element that is a loop in the dual $M^{*}$ of the matroid $M$. Equivalently, a coloop is an element that is in every base of $M$. For a cycle matroid $M_{G}$, loops correspond to loop edges and coloops correspond to bridges in the graph $G$.
Hence it was suggested that the inequalities
$\max(T_{M}(2,0),T_{M}(0,2)\geqslant T_{M}(1,1),$
$T_{M}(2,0)+T_{M}(0,2)\geqslant 2T_{M}(1,1),$
$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}$
may hold true for all matroids $M$ without loops and coloops. (These versions appear explicitly in [10], but were treated much earlier without explicitly calling them conjectures.) Note that for general matroids, all these versions are equivalent in the following sense: if one of them is true for all matroids, then the others are also true for all matroids. Applying the maximum version to $M\oplus M^{*}$ with $M^{*}$ being the dual of $M$ leads to the multiplicative version of the conjecture. (Here $M\oplus N$ denotes the direct sum of the matroids $M$ and $N$.)
2.2. Permutation Tutte polynomial
The proof of Theorem 1.3 heavily relies on the theory of permutation Tutte polynomial. The idea is that Tutte polynomial $T_{G}(x,y)$ can be written as a sum of permutation Tutte polynomials $\widetilde{T}_{H_{j}}(x,y)$ for certain bipartite graphs $H_{j}$. As a consequence certain inequalities valid for the permutation Tutte polynomial transfers to the Tutte polynomial.
Definition 2.1 ([1]).
Let $H=(A,B,E)$ be a bipartite graph. Suppose that $V(H)=[m]$. For a permutation $\pi:[m]\to[m]$, we say that a vertex $i\in A$ is internally active if
$\pi(i)>\max_{j\in N_{H}(i)}\pi(j),$
where the maximum over an empty set is set to be $\infty$- ∞. Similarly, we say that vertex $j\in B$ is externally active if
$\pi(j)>\max_{i\in N_{H}(j)}\pi(i).$
Let $\mathrm{ia}(\pi)$ and $\mathrm{ea}(\pi)$ be the number of internally and externally active vertices in $A$ and $B$, respectively. Let
$\widetilde{T}_{H}(x,y)=\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}(\pi)}y^{\mathrm{ea}(\pi)}.$
We will call $\widetilde{T}_{H}(x,y)$ the permutation Tutte polynomial of $H$.
The above definition is motivated by the following theorem of Tutte.
Theorem 2.2 (Tutte [22]).
Let $G$ be a connected graph with $m$ edges. Label the edges with $1,2,\dots,m$ arbitrarily. In the case of a spanning tree $T$ of $G$, let us call an edge $e\in E(T)$ internally active if $e$ has the largest label among the edges in the cut determined by $T$ and $e$ by removing $e$ from $T$. Let us call an edge $e\notin E(T)$ externally active if $e$ has the largest label among the edges in the cycle determined by $T$ and $e$ by adding $e$ to $T$. Let $\mathrm{ia}(T)$ and $\mathrm{ea}(T)$ be the number of internally and externally active edges, respectively. Then
$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}(T)}y^{\mathrm{ea}(T)},$
where the summation goes for all spanning trees of $G$.
Theorem 2.2 was originally a definition for the Tutte polynomial [22]. This characterization of the Tutte polynomial immediately shows that the coefficients of the Tutte polynomial are non-negative. In this theorem, we are restricted to the same labelling of the edges for all spanning trees. For those who have never seen this definition before, it might be very surprising that the Tutte polynomial is independent of the actual choice of the labelling.
To explain the connection between $T_{G}(x,y)$ and $\widetilde{T}_{H}(x,y)$, we need the concept of the local basis exchange graph.
Definition 2.3.
The local basis exchange graph $H[T]$ of a graph $G=(V,E)$ with respect to a spanning tree $T$ is defined as follows. The graph $H[T]$ is a bipartite graph whose vertices are the edges of $G$. One bipartite class consists of the edges of $T$, the other consists of the edges of $E\setminus T$, and we connect a spanning tree edge $e$ with a non-edge $f$ if $f$ is in the cut determined by $e$ and $T$, equivalently, $e$ is in the cycle determined by $f$ and $T$.
Clearly, this definition works for general matroids and their basis: if $A$ is a basis, then $e\in A$ and $f\in E\setminus A$ are adjacent in the bipartite graph $H[A]$ if $A-e+f$ is again a basis.
Figure 1 depicts a graph $G$ with a spanning tree $T$ and the bipartite graph $H[T]$ obtained from $T$.
For a fixed labelling of the edges of $G$, we get a labelling of the vertices of $H[T]$, and the internally (externally) active edges of $G$ correspond to internally (externally) active vertices of $H[T]$, so the two definitions of internal and external activity are compatible. The following lemma is crucial for us, so we even included its proof.
11112222333344446666555577778888
11112222333355557777444466668888
Figure 1. Example for a graph $G$ and the local basis exchange graph $H[T]$ obtained from a spanning tree $T$.
Lemma 2.4 (Beke, Csáji, Csikvári, Pituk [1]).
Let $G$ be a graph. For each spanning tree $T$ of $G$, let $H[T]$ be the local basis exchange graph with respect to $T$. Then
$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y),$
where the sum is over the set of spanning trees $\mathcal{T}(G)$ of $G$.
Proof.
For a fixed spanning tree $T$ and a permutation $\pi$ of the edges, the internally and externally active edges correspond to the internally and externally active vertices of $H[T]$. Hence
$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}.$
Now averaging it for all permutations $\pi\in S_{m}$ we get that
$T_{G}(x,y)$ $=\frac{1}{m!}\sum_{\pi\in S_{m}}T_{G}(x,y)$
$=\frac{1}{m!}\sum_{\pi\in S_{m}}\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
$=\sum_{T\in\mathcal{T}(G)}\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
$=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y).$
Remark 2.5.
The local basis exchange graph $H[T]$ has an isolated vertex if and only if $G$ contains a bridge or a loop. Furthermore, $H[T]$ is connected if and only $G$ is $2$-connected.
The following lemma enables us to study quadratic inequalities of the Tutte polynomial.
Lemma 2.6 (Transfer lemma [1]).
Let $x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}\geq 0$. Suppose that for any bipartite graph $H$, we have
$\widetilde{T}_{H}(x_{1},y_{1})\widetilde{T}_{H}(x_{2},y_{2})\geq\widetilde{T}_{H}(x_{0},y_{0})^{2}.$
Then for any graph $G$, we have
$T_{G}(x_{1},y_{1})T_{G}(x_{2},y_{2})\geq T_{G}(x_{0},y_{0})^{2}.$
Remark 2.7.
For us the following version of the transfer lemma will be a bit more convenient. If for any basis $A$ of a matroid $M$, the local basis exchange graph $H[A]$ satisfies that
$\widetilde{T}_{H[A]}(x_{1},y_{1})\widetilde{T}_{H[A]}(x_{2},y_{2})\geq\widetilde{T}_{H[A]}(x_{0},y_{0})^{2},$
then
$T_{M}(x_{1},y_{1})T_{M}(x_{2},y_{2})\geq T_{M}(x_{0},y_{0})^{2}.$
The proof of this claim is exactly the same as of the transfer lemma.
A key example for bounding the permutation Tutte polynomial is the following theorem proved in [1].
Theorem 2.8 ([1]).
Let $H$ be an arbitrary bipartite graph, and let $d_{i}$ be the degree of vertex $i$. Suppose that $0\leq x\leq 1$ and $y\geq 1$ or $0\leq y\leq 1$ and $x\geq 1$. Then
$\widetilde{T}_{H}(x,y)\geq\prod_{i\in A}\left(1+\frac{x-1}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(1+\frac{y-1}{d_{j}+1}\right).$
The proof of Lemma 2.8 relies on following inequality of Harris that is also a special case of the FKG-inequality [11].
Lemma 2.9 (Harris [12], Fortuin, Kasteleyn, Ginibre [11]).
Suppose that $\mu=\mu_{1}\otimes\dots\otimes\mu_{N}$ is a product measure on a set $S_{1}\times S_{2}\times\dots\times S_{N}$, and $X_{1},\dots,X_{t}$ are non-negative monotone increasing functions in the sense that if $x_{i}\geq x_{i}^{\prime}$ for $i=1,\dots,N$, then for $1\leq j\leq t$ we have
$X_{j}(x_{1},\dots,x_{N})\geq X_{j}(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
Then
$\mathbb{E}_{\mu}\left[\prod_{j=1}^{t}X_{j}\right]\geq\prod_{j=1}^{t}\mathbb{E}_{\mu}[X_{j}].$
Furthermore, if $X$ is monotone increasing and $Y$ is monotone decreasing, then
$\mathbb{E}[XY]\leq\mathbb{E}[X]\mathbb{E}[Y].$
In this paper, we will use the Harris inequality for both $[0,1]^{N}$ and the set $\{0,1\}^{A}$ for some set $A$.
In what follows, we repeatedly use the following crucial idea to express $\widetilde{T}_{H}(x,y)$. We can create a random ordering of the vertices of $H$ as follows: for each vertex $i$ we choose a uniform random number $x_{i}$ from the interval $[0,1]$. The numbers $x_{i}$ then determine an ordering of the edges. The probability that two numbers are equal is 00.
Lemma 2.10 ([1]).
Let $H$ be a bipartite graph and let $\widetilde{T}_{H}(x,y)=\sum t_{i,j}(H)x^{i}y^{j}$. Let $v(H)=m$ and let $x_{1},x_{2},\dots x_{m}$ be i.i.d. random variables with distribution $x_{i}\sim U(0,1)$. Let $I(A)=\left|\left\{v\in A|\ x_{v}\geq x_{v^{\prime}}\text{ for }v^{\prime}\in N_{H}(v)\right\}\right|$ and $I(B)=\left|\left\{v\in B|\ x_{v}\geq x_{v^{\prime}}\text{ for }v^{\prime}\in N_{H}(v)\right\}\right|$. Then
$\mathbb{P}\left(I(A)=i,I(B)=j\right)=t_{i,j}(H).$
In what follows we do a little trick. For $i\in A$ we generate $x_{i}\sim U(0,1)$ as before, but for $j\in B$ we actually first generate a uniformly random number $y_{j}$ from $[0,1]$ and let $x_{j}=1-y_{j}$. The role of this trick will be apparent soon.
For $i\in A$, let us introduce the random variable
$X_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$
and for $j\in B$, let
$Y_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
Lemma 2.11.
(a) We have
$\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{i\in A}X_{i}\cdot\prod_{j\in B}Y_{j}\right].$
(b1) If $x\geq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function for each $i\in A$.
(b2) If $0\leq x\leq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function for each $i\in A$.
(b3) For $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
(b4) Finally, for $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing for each $j\in B$.
Remark 2.12.
This lemma is the reason why we generated $x_{j}$ by $x_{j}=1-y_{j}$ for $j\in B$. If we consider the function
$X^{\prime}_{i}(x_{i},\{x_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}>x_{i}.\end{array}\right.$
instead of $X_{i}(x_{i},\{y_{j}\}_{j\in B})$, this would be neither increasing, nor decreasing.
We will also need a very little extension of the above ideas, where each vertex gets its own activity: let For $i\in A$, let us introduce the random variable
$\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x^{(i)}&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$
and for $j\in B$, let
$\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y^{(j)}&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
The following lemma is just a trivial extension of the previous lemma together with Harris’ inequality.
Lemma 2.13.
(a1) If $i\in A$ and $x^{(i)}\geq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function.
(a2) If $i\in A$ and $0\leq x^{(i)}\leq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function.
(a3) For $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
(a4) Finally, if $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing.
(b) If $x^{(i)}\geq 1$ for all $i\in A$ and $0\leq y^{(j)}\leq 1$, then
$\mathbb{E}\left[\prod_{i\in A}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]\geq\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]=\prod_{i\in A}\left(\frac{x^{(i)}}{d_{i}+1}+\frac{d_{i}}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(\frac{y^{(j)}}{d_{j}+1}+\frac{d_{j}}{d_{j}+1}\right).$
We will use one more lemma from the paper [1], namely the gluing lemma. Originally, this lemma used the condition that $H_{1}$ and $H_{2}$ are trees, but the proof never used this condition.
Lemma 2.14 (Gluing lemma [1]).
Let $x\geq 1$ and $0\leq y\leq 1$. Let $H_{1}$ be a rooted graph with root vertex $v_{1}$. Let $H_{2}$ be another rooted graph with root vertex $v_{2}$. Let $H$ be obtained from $H_{1}$ and $H_{2}$ by identifying $v_{1}$ and $v_{2}$ in the union of $H_{1}$ and $H_{2}$. Let $v$ be the vertex obtained from identifying $v_{1}$ and $v_{2}$. Assume that the bipartite parts of $H$ determines the bipartite parts of $H_{1}$ and $H_{2}$, that is, if $v\in A(H)$, then $v_{1}\in A(H_{1})$ and $v_{2}\in A(H_{2})$, and if $v\in B(H)$, then $v_{1}\in B(H_{1})$ and $v_{2}\in B(H_{2})$.
(a) If $v\in A$, then
$x\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
(b) If $v\in B$, then
$\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
(c) In particular,
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\frac{1}{x}(\widetilde{T}_{H_{1}}(x,0)\widetilde{T}_{H_{1}}(0,x))(\widetilde{T}_{H_{2}}(x,0)\widetilde{T}_{H_{2}}(0,x))$
3. Warm-up: asymptotic computation of some permutation Tutte polynomial
In this section we motivate an important technique of this paper by computing the asymptotic value of the permutation Tutte polynomial of certain bipartite graphs.
Let us start with a very simple example, the complete bipartite graph.
Theorem 3.1.
We have
[…]
Proof.
We generate the random permutation on $V(H)=A\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let $t=\max_{v\in A}x_{v}$ and $s=\max_{v\in B}x_{v}$. We have $ab$ choices for the vertices that take the value $t$ and $s$, let these vertices be $v_{A}$ and $v_{B}$. If $s>t$, then the permutation has weight 00 as $v_{B}$ is an active vertex. If $t>s$, then no vertex can be active in $B$. For each $w\in A$ that is not equal to $v_{A}$ two things can happen: if $x_{w}<s$, then $w$ is not active, so $X_{w}=1$, or $s<x_{w}<t$ and then $w$ is active so $X_{w}=x$. For a vertex $w\in B$ we simply need to have $x_{w}<s$ and $w$ will not be active. The integral formula then follows. ∎
Remark 3.2.
Suppose that $a=\alpha m$ and $b=\beta m$, where $\alpha,\beta$ are fixed such that $\alpha+\beta=1$ and $m\to\infty$. Then the exponential growth constant of $\widetilde{T}_{K_{a,b}}(x,0)$ is simply
$\lim_{m\to\infty}\widetilde{T}_{K_{a,b}}(x,0)^{1/m}=\max_{s,t}s^{\beta}(s+x(t-s))^{\alpha}.$
Clearly, at the maximum we have $t=1$ and we simply need to maximize $s^{\beta}(s+x(1-s))^{\alpha}$. This turns out to be at $s=\min\left(1,\frac{\beta x}{\beta-1}\right)$. If $\beta<\frac{x-1}{x}$, then the growth constant is
$\left(\frac{\beta x}{\beta-1}\right)^{\beta}\left(x+(1-x)\frac{\beta x}{x-1}\right)^{\alpha}=\alpha^{\alpha}(1-\alpha)^{1-\alpha}\frac{x}{(x-1)^{1-\alpha}}.$
If $\beta\geq\frac{x-1}{x}$, then the exponential growth constant is simply $1$.
Let us consider the graph $H_{a,b,c}$ introduced in [1]: we start with a complete bipartite graph $K_{a,b}$ with vertex set $A\cup B$, and then attach $c$ pendant leaves to $c$ distinct vertices of $B$, let $C$ be the set of these leaf vertices. So the resulting bipartite graph has $a+c$ vertices on one side and $b$ vertices on the other side. The graphs $H_{n,n,n}$ played an important role in the refutation of the matroidal version of the Merino–Welsh conjecture.
Figure 2. The graph $H_{6,6,6}$.
It turns out that for even $n$ there are matroids for which all local basis exchange graphs are isomorphic to $H_{n,n,n}$. Indeed, all local basis exchange graphs of the matroid $U^{(2)}_{\frac{3}{2}n,n}$ are isomorphic to $H_{n,n,n}$, where $U_{\frac{3}{2}n,n}$ is the uniform matroid of rank $n$ on $\frac{3}{2}n$ elements, and the matroid $M^{(2)}$ is obtained from $M$ by replacing each element of $M$ with two parallel elements. Furthermore, we have
$\widetilde{T}_{H_{n,n,n}}(2,0)\widetilde{T}_{H_{n,n,n}}(2,0)<\widetilde{T}_{H_{n,n,n}}(1,1)^{2}$
for large enough $n$.
We can now give a simple proof of this fact.
Theorem 3.3.
We have
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}=\max_{s\in[0,1]}\left(s+x(1-s)\right)\left(xs+(1-x)\frac{s^{2}}{2}\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}&\text{if}\ \ x\geq\sqrt{3},\\ \frac{1}{2}(x+1)&\text{if}\ \ 1<x\leq\sqrt{3}\end{cases}$
and
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}=\max_{s\in[0,1]}t\left(\frac{t^{2}}{2}+\left(\frac{1}{2}-\frac{t^{2}}{2}\right)x\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}&\text{if}\ \ x\geq\frac{3}{2},\\ \frac{1}{2}&\text{if}\ \ 1<x\leq\frac{3}{2}.\end{cases}$
In particular, if $x\geq\sqrt{3}$ we have
$\lim_{n\to\infty}\left(\widetilde{T}_{H_{n,n,n}}(x,0)\widetilde{T}_{H_{n,n,n}}(0,x)\right)^{1/n}=\left(\frac{x^{3}}{9(x-1)}\right)^{3/2}.$
Sketch of the proof..
As before we generate the random permutation on $V(H)$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let
$t=\max_{v\in A}x_{v}\ \ \ \text{and}\ \ \ s=\max_{v\in B}x_{v}.$
Let $v_{A}$ and $v_{B}$ be the vertices, where these maximum are achieved. (Note that $t$ is only maximum on $A$ and not on $A\cup C$.) If $s<t$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
[…]
Here $s+x(t-s)$ is the contribution of a vertex $w\in A$ that is not $v_{A}$. If $x_{w}=s_{i}<s$ for some $w\in B$ that is not $v_{B}$, then the attached leaf in $C$ has contribution $s_{i}+x(1-s_{i})$. (Note that attached leaf vertex $u$ can have a value bigger than $t$ as $t$ was only a maximum among the vertices of $A$.) The last term $s+x(1-s)$ is simply the contribution of the leaf attached to $v_{B}$.
If $t<s$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
[…]
Here the contribution of a vertex $w\in A$ is simply $t$ as it can never be active since $s<t$. If we have a vertex $w\in B$ that is not $v_{B}$ with $x_{w}=s_{i}<s$, then we need to distinguish two cases. If $s_{i}<t$, then $w$ is not active, and so the attached leaf has a contribution $s_{i}+x(1-s_{i})$. If $s_{i}>t$ then, in order to make $w$ inactive the attached leaf $u$ should have a value $x_{u}>s_{i}$, but then it is automatically active. The contribution of the leaf attached to $v_{B}$ is $x(1-s)$ as it has to be active to make $v_{B}$ inactive.
We have
$\lim_{n\to\infty}I_{1}^{1/n}$ $=\max_{0\leq s\leq t\leq 1}(s+x(t-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
$=\max_{0\leq s\leq 1}(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
The function $(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$ has either maximum at $\left(1-\frac{1}{\sqrt{3}}\right)\frac{x}{x-1}$ if this is less than $1$, or at $s=1$. In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}$. In the second case, its value is $\frac{1}{2}(x+1)$.
On the other hand,
$\lim_{n\to\infty}I_{2}^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left((1-x)\frac{t^{2}}{2}+xt+x(s-t)-x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
$=\max_{0\leq t\leq s\leq 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
$=\max_{0\leq s\leq 1}\left(x\left(s-\frac{s^{2}}{2}\right)+\frac{1}{2}\right)$
$=\frac{1}{2}(x+1),$
so this is always at most $\lim_{n\to\infty}I_{1}^{1/n}$.
Then we have
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}$ $=\lim_{n\to\infty}(I_{1}+I_{2})^{1/n}$
$=\max\left(\lim_{n\to\infty}I_{1}^{1/n},\lim_{n\to\infty}I_{2}^{1/n}\right)$
$=\lim_{n\to\infty}I_{1}^{1/n}.$
This proves the first statement. The second statement is actually simpler. We have
[…]
From this we get that
$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
$=\max_{0\leq t\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{1}{2}-\frac{t^{2}}{2}\right)\right)$
This expression has either maximum at $x=\left(\frac{x}{3(x-1)}\right)^{1/2}$ if $x\geq\frac{3}{2}$ or at $1$ if $x\leq\frac{3}{2}$. In the first case its value is $\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}$. In the second case, its value is $\frac{1}{2}$. This completes the sketch of the proof. ∎
Remark 3.4.
It is clear from the above examples that even though the permutation Tutte polynomial might be complicated still it is possible to compute the exponential growth of it for many families of bipartite graphs. Another important take-away that it is often possible to understand those permutations that contributes the most weight to the permutation Tutte polynomial, and concentrating for these permutations might give a better lower or upper bound for the permutation Tutte polynomial.
4. The main lemma
In this section we give the main ingredient of the proofs of Theorems 1.3 and 1.4.
Theorem 4.1.
Let $H=(A,B,E)$ be a bipartite graph. For $x\geq 2$ and a fixed $0\leq s\leq 1$ let us introduce the function
$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}.$
Then
$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
Proof.
Recall that
$\widetilde{T}_{H}(x,0)=\int_{[0,1]^{A}}\int_{[0,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
The first idea is that we only take those values for which $x_{u}=1-y_{u}\leq s$ for $u\in B$. Hence
$\widetilde{T}_{H}(x,0)\geq\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
For each $v\in A$ we decompose the integral over $[0,1]$ to $[0,s]$ and $[s,1]$, this way we decomposed the integral $[0,1]^{A}$ to the sum of $2^{|A|}$ integrals. Since $x_{u}\leq s$ for $u\in B$ we immediately get that if $x_{v}\geq s$, then $X_{v}(\underline{x},\underline{y})=x$. Let $T\subseteq A$ be the set of those vertices $v\in A$ for which $x_{v}\geq s$. Then
$\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
$=\sum_{T\subseteq A}(x(1-s))^{|T|}\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
Let $N(T)\subseteq B$ be the neighbors of $T$. If $u\in N(T)$, then $\widehat{Y}_{u}(\underline{x},\underline{y}):=Y_{u}(\underline{x},\underline{y})=1$, because the vertex $u$ cannot be active anymore. Hence
$I_{T}:=\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
corresponds to $s^{|A\setminus T|+|B|}$ times a permutation Tutte polynomial, where an active vertex $i\in A\setminus T$ gets a weight $x^{(i)}=x$ while in $B$ an active vertex $j$ gets a weight $y^{(j)}=1$ if it is $N(T)$, and $y^{(j)}=0$ if it is $B\setminus N(T)$, so for this term we have a lower bound from Lemma 2.13:
$I_{T}$ $=s^{|A\setminus T|+|B|}\mathbb{E}\left[\prod_{i\in A\setminus T}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]$
$\geq s^{|A\setminus T|+|B|}\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]$
$=s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
So far we got that
$\widetilde{T}_{H}(x,0)$ $=\sum_{T\subseteq A}(x(1-s))^{|T|}I_{T}$
$\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
(4.1) $=\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
Next let
$Z=\prod_{v\in A}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)\right),$
and let us consider the measure $\mu$ on $\{0,1\}^{A}$ for which
$\mu(T)=\frac{1}{Z}(x(1-s))^{|T|}\prod_{A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right).$
Note that $\mu$ is a product measure $\mu=\bigoplus_{v\in A}\mu_{v}$, where
$\mu_{v}(0)=\frac{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s}{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)}\ \ \ \text{and}\ \ \ \mu_{v}(1)=\frac{x(1-s)}{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)}.$
For each $u\in B$ consider the function
$f_{u}(T)=\begin{cases}s&\text{if }u\in N(T),\\ \frac{d_{u}}{d_{u}+1}s&\text{if }u\not\in N(T).\end{cases}$
Then we can rewrite inequality (4.1) as follows:
$\widetilde{T}_{H}(x,0)$ $\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
$=Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T).$
All functions $f_{u}$ are monotone increasing on the space $\{0,1\}^{A}$ identified with the subsets $T$ of $A$, since adding vertices to $T$ can only increase $N(T)$, so by Harris’ inequality we have
$Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T)\geq Z\prod_{u\in B}\left(\sum_{T\subseteq A}\mu(T)f_{u}(T)\right).$
Here
$\sum_{T\subseteq A}\mu(T)f_{u}(T)=\sum_{T\subseteq A\atop u\in N(T)}\mu(T)s+\sum_{T\subseteq A\atop u\not\in N(T)}\mu(T)s\left(1-\frac{1}{d_{u}+1}\right)$
$=s-\frac{s}{d_{u}+1}\sum_{T\subseteq A\atop u\not\in N(T)}\mu(T)=s-\frac{s}{d_{u}+1}\sum_{T\subseteq A\atop T\cap N(u)=\emptyset}\mu(T)$
$=s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\frac{\frac{d_{v}+x}{d_{v}+1}s}{x(1-s)+\frac{d_{v}+x}{d_{v}+1}s}=s\left(1-\frac{1}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
Hence
$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
By applying Theorem 4.1 for both $\widetilde{T}_{H}(x,0)$ and $\widetilde{T}_{H}(0,x)$ we get the following corollary.
Corollary 4.2.
We have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right].$
Lemma 4.3.
If $x\geq 1$ and $0\leq s\leq 1$, then the function $\gamma_{x,s}(d)$ is monotone decreasing in $d$. Furthermore, $\gamma_{x,s}(d)\leq s$.
Proof.
The function $\gamma_{x,s}(d)$ is a rational function in $d$:
$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}=\frac{sd+sx}{(x(1-s)+s)d+x}$
In general, the derivative of a function $f(t)=\frac{at+b}{ct+d}$ is $\frac{ad-bc}{(ct+d)^{2}}$, so it is enough to check that
$sx-sx(x(1-s)+s)=sx(1-x)(1-s)\leq 0$
which is clearly true. The second statement follows from $d+x\leq(d+1)x$ as follows:
$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}\leq\frac{(d+x)s}{(d+x)s+(d+x)(1-s)}=s.$
Definition 4.4.
Let
$G(d,x,s,\gamma):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right).$
For any $10\leq\gamma<$ 1 we have
$G(\infty,x,s):=\lim_{d\to\infty}G(d,x,s,\gamma)=(s+x(1-s))s.$
By combining Corollary 4.2 with Lemma 4.6 we immediately get the following statement.
Corollary 4.5.
Let $H$ be a bipartite graph with minimum degree at least $\delta$. Then for $x\geq 2$ and $0\leq s\leq 1$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geq\prod_{v\in V}G(d_{v},x,s,s).$
Proof.
We have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right]$
$\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(\delta)\right)\right]$
$=\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))$
$\geq\prod_{v\in V}G(d_{v},x,s,s).$
Lemma 4.6.
Let $x\geq 2$ and $\gamma\leq s<1$. A sufficient condition for
$G(d,x,s,\gamma)\geq G(d+1,x,s,\gamma)$
is $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq\gamma.$
Furthermore, the function $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}$ is monotone increasing in $d$ for $x\geq 1$ and $d\geq 2$, so it is enough to check the above inequality for some fixed $d_{0}$ to conclude that the sequence $G(d,x,s,\gamma)$ is decreasing for $d\geq d_{0}$.
Proof.
We have
$G(d,x,s,\gamma)-G(d+1,x,s,\gamma)$
$=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right)-\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)\left(s-\frac{s}{d+2}\gamma^{d+1}\right)$
$=s^{2}\left(\frac{d+x}{d+1}-\frac{d+1+x}{d+2}\right)-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)+\frac{s}{d+2}\gamma^{d+1}\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)$
$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}\gamma^{d}$
which is non-negative if $\frac{x-1}{x(d+2)}\geq\gamma^{d-1}$. Finally, the derivative of $h(t):=\frac{1}{t-1}\ln\left(\frac{x-1}{x(t+2)}\right)$ is
$h^{\prime}(t)=\frac{1}{(t-1)^{2}}\ln\left(\frac{x(t+2)}{x-1}\right)-\frac{1}{(t-1)(t+2)}.$
Then
$h^{\prime}(t)\geq\frac{\ln(t+2)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}\geq\frac{\ln(3)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}>0$
if $t>1$. ∎
5. Proof of Theorem 1.3.
In this section we prove that
$T_{M}(x,0)T_{M}(0,x)\geq T_{M}(1,1)^{2}$
whenever $x\geq 2.355$. By the transfer lemma it is enough to prove that
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
for any bipartite graph $H$ with minimum degree at least $1$. Note that $\widetilde{T}_{H}(1,1)=1$.
We do not immediately prove the best value $2.355$, but give various improvements to the value $2.9243$ step by step to digest one idea at a time. All computations are carried out in the jupyter notebook [7] and we also give the tables in the Appendix.
Idea 1. By Corollary 4.5 we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(1)).$
Since $\widetilde{T}_{H}(1,1)^{2}=1$ it is enough if $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for all $d\geq 1$. Let
$x=\frac{3+\sqrt{5}}{2}\approx 2.6180..\ \ \text{and}\ \ \ s=\frac{6+2\sqrt{5}}{7+3\sqrt{5}}\approx 0.7639...$
These numbers are chosen in such a way that $G(1,x,s,\gamma_{x,s}(1))=1$. It is easy to check that for $d=9$ we have $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq s\geq\gamma_{x,s}(1)$. By Lemma 4.6 this implies that for $d\geq 9$ we have $G(d,x,s,\gamma_{x,s}(1))\geq G(\infty,x,s)>1$. It is also easy to check (see the jupyter notebook or Table 1 in the Appendix) that $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for $d\in\{2,\dots,10\}$. So with $x=\frac{3+\sqrt{5}}{2}$ we have
$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
implying that
$T_{M}(x,0)T_{M}(0,x)\geq T_{M}(1,1)^{2}$
for all matroids without loops and coloops, in particular for all graphs $G$ without loops and bridges.
Idea 2. Observe that we can assume that $H$ is connected since $\widetilde{T}_{H}(x,y)=\prod_{i=1}^{k}\widetilde{T}_{H_{i}}(x,y)$ if $H$ has connected components $H_{1},\dots,H_{k}$. The statement $\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq 1$ is trivially true for $H=K_{2}$, the complete graph on $2$ vertices. So we can assume that $H$ has at least $3$ vertices. On the other hand, if $H$ is connected and has at least $3$ vertices, then a degree one vertex cannot be adjacent to another degree one vertex. Thus it is enough if
$G(1,x,s,\gamma_{x,s}(2))\geq 1\ \ \text{and}\ \ G(d,x,s,\gamma_{x,s}(1))\geq 1$
for all $d\geq 2$. This is satisfied if $x=2.54$ and $s=0.76$. (See Table 2.)
Idea 3. Observe that it is not really necessary that $G(1,x,s,\gamma)\geq 1$, it is quite enough if for a non-leaf vertex the product with its leaf neighbors is at least $1$. If $x\geq 2.2$, then any vertex can have at most $2$ leaf neighbors, since for the star $S_{4}$ on $4$ vertices we have
$\frac{1}{x}\widetilde{T}_{S_{4}}(x,0)\widetilde{T}_{S_{4}}(0,x)=\frac{1}{x}\cdot\frac{x^{3}+x^{2}+x}{4}\cdot\frac{x}{4}=\frac{x^{3}+x^{2}+x}{16}>1$
so by the gluing lemma (Lemma 2.14) we know that a graph containing a pending $S_{4}$ cannot be a minimal counter example. So it is enough to check if for every $d\geq 2$ we have
$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$
and
$G(d,x,s,\gamma_{x,s}(1))\geq 1.$
In order to check the first inequality for large $d$ it is enough if for some $d_{0}$ we have
$G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1,$
then we get that
$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$
for $d\geq d_{0}$. So we only need to check the first inequality for $d=2,\dots,d_{0}$ and the inequality $G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1$. These inequalities are satisfied for $x=2.36$, $s=0.78$ and $d_{0}=44$. Again see the jupyter notebook [7] or alternatively, Table 3.
Idea 4. We can gain a very small further improvement as follows. For $d\geq 2$ we can use that
$\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)$
$\geq\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\gamma_{x,s}(1)^{\min(2,d_{v}-1)}\gamma_{x,s}(2)^{d_{v}-\min(2,d_{v}-1)}\right),$
Let
$G_{2}(d,x,s,\gamma_{1},\gamma_{2}):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma_{1}^{\min(2,d-1)}\gamma_{2}^{d-\min(2,d-1)}\right).$
Then it is enough to check that
$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\geq 1$
for $d\geq 2$ and
$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))G(d,x,s,\gamma_{x,s}(1))^{\min(2,d-1)}\geq 1$
for all $d\geq 2$. This leads to the tiny improvement $x=2.355$ ($s=0.78$ and $d=100$). See [7] or Table 4 in the Appendix for further details.
6. Matroids with prescribed circuit lengths
In this section, we prove Theorem 1.4. The heart of the argument is the following lemma.
Lemma 6.1.
Let $k\geq 4$ be a real number. Let $H$ be a bipartite graph such that all degrees of $H$ lie in the interval $[k+1,k^{4}-2k^{2}-1]$. Then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
Proof.
We will use that $\widetilde{T}_{H}(1,1)=1$, and Corollary 4.5:
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\prod_{v\in V}G(d_{v},2,s,s).$
Note that
$G(d,2,s,s)=\left(\frac{d+2}{d+1}\cdot s+2(1-s)\right)\left(s-\frac{s^{d+1}}{d+1}\right)=\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)\left(s-\frac{s^{d+1}}{d+1}\right).$
We will show that if $s=1-\frac{1}{k^{2}}$, then for all $d\in[k+1,k^{4}-2k^{2}-1]$ we have $G(d,2,s,s)\geq 1$. First observe that $G(d,2,s,s)$ is a concave function of $s$ on the interval $(0,1)$ as its second derivative is
$-\frac{(2-s)(d^{2}+d)s^{d}+2ds}{(d+1)s},$
which is clearly negative. This means that $I_{d}:=\{s\ |\ G(d,2,s,s)\geq 1\}$ is an interval for each $d$. We will show that $1-\frac{1}{k^{2}}\in I_{k+1}\cap I_{k^{4}-2k^{2}-1}$. By first choosing $k=d-1$ we get that $1-\frac{1}{(d-1)^{2}}\in I_{d}$. If we choose $k^{2}=1+\sqrt{d+2}$, then $k^{4}-2k^{2}-1=d$ and so $1-\frac{1}{1+\sqrt{d+2}}\in I_{d}$. Hence $\left[1-\frac{1}{1+\sqrt{d+2}},1-\frac{1}{(d-1)^{2}}\right]\subseteq I_{d}$. This means that if $k+1\leq d\leq k^{4}-2k^{2}-1$, then $1-\frac{1}{k^{2}}\in I_{d}$.
First we show that $1-\frac{1}{k^{2}}\in I_{k+1}$, that is, $G(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}})\geq 1$. We will use that
$(1-t)^{r}\leq 1-\binom{r}{1}t+\binom{r}{2}t^{2}$
for $0\leq t\leq 1$ by Bonferroni’s inequality applied to independent events with probability $t$. Hence
$G\left(d,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $=\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{1}{k^{2}}\right)^{d+1}\right)$
$\geq\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{d+1}{k^{2}}+\frac{(d+1)d}{2k^{2}}\right)\right)$
For $d=k+1$ we get that
$G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $\geq\left(\frac{k+3}{k+2}+\frac{k+1}{k+2}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{k+2}\left(1-\frac{k+2}{k^{2}}+\frac{(k+2)(k+1)}{2k^{2}}\right)\right)$
$=1+\frac{1}{2}\frac{3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2}{k^{8}+4k^{7}+4k^{6}}$
For $k\geq 4$ we have $03k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2>$ 0 showing that $G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)>1$.
Next we show that $1-\frac{1}{k^{2}}\in I_{k^{4}-2k^{2}-1}$. For $s=1-\frac{1}{k^{2}}$ and $d=k^{4}-2k^{2}-1$ we have
$\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)s=1+\frac{1}{k^{8}-2k^{6}}$
and
$\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)\frac{s^{d+1}}{d+1}=\frac{k^{6}-k^{4}-k^{2}-1}{k^{6}-2k^{4}}\cdot\frac{1}{k^{4}-2k^{2}}\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}.$
So we need to prove that
$\frac{1}{k^{8}-2k^{6}}>\frac{k^{6}-k^{4}-k^{2}-1}{k^{6}-2k^{4}}\cdot\frac{1}{k^{4}-2k^{2}}\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}$
which is equivalent with
$\frac{k^{2}-2}{k^{6}-k^{4}-k^{2}-1}>\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}.$
By $1-t<e^{-t}$ we have
$\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}<e^{-\frac{1}{k^{2}}(k^{4}-2k^{2})}=e^{-k^{2}-2}.$
Let us introduce $y=k^{2}-2$. Then $k^{6}-k^{4}-k^{2}-1=y^{3}+5y^{2}+7y+1$, so we need that
$\frac{y}{y^{3}+5y^{2}+7y+1}<e^{-y}$
or equivalently
$y^{2}+5y+7+\frac{1}{y}<e^{y}.$
Since $k\geq 3$ we have $y=k^{2}-2\geq 7$, but $e^{y}>y^{2}+5y+7+\frac{1}{y}$ already true for $y\geq 4$.
Remark 6.2.
There was no attempt to optimize Lemma 6.1, not even concerning the minimum degree. It turns out that if $s=0.9226$, then
$G(d,2,s,\gamma_{2,s}(3))>1$
for $3\leq d\leq 141$. So if all degrees of a bipartite graph $H$ lies in the interval $[3,141]$, then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
Similarly, for $s=0.9622$ we have
$G(d,2,s,\gamma_{2,s}(4))>1$
for $4\leq d\leq 646$. So if all degrees of a bipartite graph $H$ lies in the interval $[4,646]$, then
$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
This shows that in Theorem 1.4 we could have written $\ell\geq 4$ and probably a faster growing function than $(\ell-2)^{4}$.
Now we are ready to finish the proof of Theorem 1.4.
Proof of Theorem 1.4.
Let $A$ be a basis of the matroid $M$ on ground set $E$. Let $B=E\setminus A$. Let us consider the local basis exchange graph $H[A]$ on $A\cup B$. Observe that $v\in B$ are adjacent with exactly those elements that are in his fundamental circuit. So $\ell\leq d_{v}+1\leq(\ell-2)^{4}$. Similarly, if $u\in A$, then it is adjacent with those elements of $B$ that are in its fundamental circuits with respect to $B$ in $M^{*}$. So $\ell\leq d_{u}+1\leq(\ell-2)^{4}$.
By choosing $k=\ell-1$ we get that $k^{4}-2k^{2}-1=(\ell-1)^{4}-2(\ell-1)^{2}-1\geq(\ell-2)^{4}$ if $\ell\geq 3$. So by Lemma 1.4 we have
$\widetilde{T}_{H[A]}(2,0)\widetilde{T}_{H[A]}(0,2)\geq\widetilde{T}_{H[A]}(1,1)^{2}.$
Since this is true for all basis $A$ of the matroid $M$, we get by the Transfer lemma (Lemma 2.6 and Remark 2.7) that
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
7. Concluding remarks
We end this paper with some conjectures.
Conjecture 7.1.
If the bipartite graph $H$ has minimum degree at least $2$, then
$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geq\widetilde{T}_{H}(1,1)^{2}.$
If this conjecture is true, then the following conjecture would be an immediate corollary.
Conjecture 7.2.
If $M$ is a simple co-simple matroid, then
$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
Acknowledgment. The author is very grateful to Csongor Beke, Gergely Kál Csáji and Sára Pituk for discussions on the topic of this paper.

### v2

@@
 Abstract.
-The Merino-Welsh conjecture states that for a graph $G$ without loops and bridges we have
-$\max(T_{G}(2,0),T_{G}(0,2))\geq T_{G}(1,1).$
-Later Jackson proved that for any matroid $M$ without loop and coloop we have
-$T_{M}(3,0)T_{M}(0,3)\geq T_{M}(1,1)^{2}.$
-The value $3$ in this statement was improved to $2.9242$ by Beke, Csáji, Csikvári and Pituk. In this paper, we further improve on this result by showing that
-$T_{M}(2.355,0)T_{M}(0,2.355)\geq T_{M}(1,1)^{2}.$
-We also prove that the Merino–Welsh conjecture is true for matroids $M$, where all circuits of $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$ for some $\ell\geq 6$.
-The research was supported by the MTA-Rényi Counting in Sparse Graphs ”Momentum” Research Group, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
+The Merino-Welsh conjecture states that for a graph $G$ without loops and bridges the Tutte polynomial $T_{G}(x,y)$ satisfies the inequality
+$\max(T_{G}(2,0),T_{G}(0,2))\geqslant T_{G}(1,1).$
+Later Jackson proved that for any matroid $M$ without loops and coloops we have
+$T_{M}(3,0)T_{M}(0,3)\geqslant T_{M}(1,1)^{2}.$
+The value $3$ in this statement was improved to $2.9243$ by Beke, Csáji, Csikvári and Pituk. In this paper, we further improve on this result by showing that
+$T_{M}(2.355,0)T_{M}(0,2.355)\geqslant T_{M}(1,1)^{2}.$
+We also prove that the Merino–Welsh conjecture is true for matroids $M$, where all circuits of $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{2}(\ell^{2}-4\ell+2)$ for some $\ell\geqslant 4$.
+The research was supported by the MTA-Rényi Counting in Sparse Graphs ”Momentum” Research Group, by the Hungarian National Research, Development and Innovation Office, Advanced grant 153378, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
 1. Introduction
 Let $G$ be a connected graph without loops and bridges. Merino and Welsh [18] conjectured that
-$\max\left(\alpha(G),\alpha^{*}(G)\right)\geq\tau(G),$
+$\max\left(\alpha(G),\alpha^{*}(G)\right)\geqslant\tau(G),$
 where $\alpha(G),\alpha^{*}(G),\tau(G)$ denote the number of acyclic orientations, strongly connected orientations, and spanning trees of $G$, respectively.
 These quantities are evaluations of the Tutte polynomial $T_{G}(x,y)$, namely $T_{G}(2,0)=\alpha(G)$, $T_{G}(0,2)=\alpha^{*}(G)$ and $T_{G}(1,1)=\tau(G)$.
-Conde and Merino [5] proposed "additive" and "multiplicative" versions of this conjecture:
-$T_{G}(2,0)+T_{G}(0,2)$ $\geq 2T_{G}(1,1)$
-$T_{G}(2,0)T_{G}(0,2)$ $\geq T_{G}(1,1)^{2},$
+Conde and Merino [5] proposed “additive” and “multiplicative” versions of this conjecture:
+$T_{G}(2,0)+T_{G}(0,2)$ $\geqslant 2T_{G}(1,1),$
+$T_{G}(2,0)T_{G}(0,2)$ $\geqslant T_{G}(1,1)^{2},$
 respectively.
@@
 Furthermore, let $x_{0}$ be the largest root of the polynomial $x^{3}-9(x-1)$.
-($x_{0}\approx 2.22668...$) Then for $0\leqslant a<x_{0}$ there are infinitely many matroids $M$ without loops and coloops for which
+($x_{0}\approx 2.22668$...) Then for any $0\leqslant a<x_{0}$ there are infinitely many matroids $M$ without loops and coloops for which
 $T_{M}(a,0)T_{M}(0,a)<T_{M}(1,1)^{2}.$
 This paper investigates the following question: for what values of a does the inequality
-$T_{M}(a,0)T_{M}(0,a)\geq T_{M}(1,1)^{2}$
+$T_{M}(a,0)T_{M}(0,a)\geqslant T_{M}(1,1)^{2}$
 hold for all loopless and coloopless matroids $M$? In this direction the first major result is due to Jackson.
@@
 For any matroid M without loops and coloops,
-$T_{M}(3,0)T_{M}(0,3)\geq T_{M}(1,1)^{2}.$
+$T_{M}(3,0)T_{M}(0,3)\geqslant T_{M}(1,1)^{2}.$
 Jackson’s result was improved by Beke, Csáji, Csikvári and Pituk in the paper [1]. They showed that one can write $2.9243$ instead of $3$. In this paper, we further improve on this inequality.
 Theorem 1.3.
-For any matroid M without loops and coloops and $a\geq 2.355$,
-$T_{M}(a,0)T_{M}(0,a)\geq T_{M}(1,1)^{2}.$
+For any matroid M without loops and coloops and $a\geqslant 2.355$,
+$T_{M}(a,0)T_{M}(0,a)\geqslant T_{M}(1,1)^{2}.$
 We may also study which matroid classes satisfy the product version of the Merino–Welsh conjecture.
-The following theorem is motivated by the fact that paving matroids satisfy the Merino–Welsh conjecture.
+The following theorem is motivated by the fact that paving matroids satisfy the Merino–Welsh conjecture [4, 10].
 A matroid of rank $r$ is a paving matroid if all circuits have length $r$ or $r+1$.
 Theorem 1.4.
-Suppose that there exists an $\ell\geq 6$ such that all circuits of the matroid $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{4}$. Then
-$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
-The proof of Theorem 1.3 and Theorem 1.4 are based on the theory of the permutation Tutte polynomial developed in the paper [1].
+Suppose that there exists an $\ell\geqslant 4$ such that all circuits of the matroid $M$ and its dual $M^{*}$ have length between $\ell$ and $(\ell-2)^{2}(\ell^{2}-4\ell+2)$. Then
+$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
+The proofs of Theorem 1.3 and Theorem 1.4 are based on the theory of the permutation Tutte polynomials developed in the paper [1].
 While Theorem 1.4 does not imply that paving matroids satisfy the Merino–Welsh conjecture, one can prove this fact by modifying the proof of Theorem 1.4.
-Notations.
+Notation.
 Throughout the paper $G=(V,E)$ is an arbitrary graph and $H=(A,B,E)$ is a bipartite graph.
 $K_{a,b}$ denotes the complete bipartite graph with parts of size $a$ and $b$.
-$S_{k}$ denotes the star graph on $k$ vertices.
+$S_{k}$ denotes the star graph on $k$ vertices, that is, $S_{k}=K_{1,k-1}$.
 For a vertex $v$ the degree of $v$ is denoted by $d_{v}$.
@@
 This paper is organized as follows.
-  • In the next section we introduce the basic concepts from matroid theory that we will use, and revisit the theory of permutation Tutte polynomial $\widetilde{T}_{H}(x,y)$ developed in the paper [1].
-  • In Section 3 we compute the growth constant $\lim_{n\to\infty}\widetilde{T}_{H_{n}}(x,0)^{1/n}$ for some family of bipartite graphs.
-While this section is not necessary for the proofs of Theorem 1.3 and 1.4 it provides an important intuition to study the permutation Tutte polynomial.
+  • In the next section, we introduce the basic concepts from matroid theory that we will use, and revisit the theory of the permutation Tutte polynomial $\widetilde{T}_{H}(x,y)$ developed in the paper [1].
+  • In Section 3 we compute the growth constant $\lim_{n\to\infty}\widetilde{T}_{H_{n}}(x,0)^{1/n}$ for some families of bipartite graphs.
+While this section is not necessary for the proofs of Theorem 1.3 and 1.4 it provides an important intuition concerning the permutation Tutte polynomial.
   • In Section 4 we prove a technical, but very important lemma that provides the basis of the proof of Theorems 1.3 and 1.4.
@@
   • In Section 6 we prove Theorem 1.4.
-  • In Section 7 we end the paper some conjectures.
+  • In Section 7 we end the paper with some conjectures.
   • In the Appendix one can find some tables that are used in the proof of Theorem 1.3.
 2. Preliminaries
-This section collects the necessary tools from matroid theory together, and recall some of the basic facts from the theory of permutation Tutte polynomial.
+This section collects the necessary tools from matroid theory, and recalls some of the basic facts from the theory of the permutation Tutte polynomial.
 2.1. Tutte polynomial and matroids
@@
 $T_{G}(x,y)=\sum_{A\subseteq E}(x-1)^{k(A)-k(E)}(y-1)^{k(A)+|A|-v(G)},$
-with $k(A)$ denoting the number of connected components of the graph $(V,A)$, see [22].
-There is a vast amount of literature on the properties of the Tutte polynomial and its applications, see for instance, [3, 6, 8, 23], or the book [9].
+with $k(A)$ denoting the number of connected components of the graph $(V,A)$ and $v(G)$ denotes the number of vertices of $G$, see [22].
+There is a vast literature on the properties of the Tutte polynomial and its applications, see for instance, [3, 6, 8, 23], or the book [9].
 The Tutte polynomial naturally extends to matroids.
-Recall that a matroid $M$ is a pair $(E,\mathcal{I})$ such that $\mathcal{I}\subseteq 2^{E}$, called independent sets, satisfies the axioms (i) $\emptyset\in\mathcal{I}$, (ii) if $A^{\prime}\subseteq A\in\mathcal{I}$, then $A^{\prime}\in\mathcal{I}$, and (iii) if $A,B\in\mathcal{I}$ such that $|B|<|A|$, then there exists an $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{I}$.
-Given a set $S\subseteq E$, the maximal independent subsets of $S$ all have the same cardinality, and this cardinality is called the rank of the matroid, denoted by $r(S)$.
-The maximum size independent sets of $M$ are called bases, and their set is denoted by $\mathcal{B}(M)$.
+Recall that a matroid $M$ is a pair $(E,\mathcal{I})$ such that $\mathcal{I}\subseteq 2^{E}$, the collection of independent sets, satisfies the axioms
+  • (i)
+    $\emptyset\in\mathcal{I}$,
+  • (ii)
+    if $A^{\prime}\subseteq A\in\mathcal{I}$, then $A^{\prime}\in\mathcal{I}$,
+  • (iii)
+    if $A,B\in\mathcal{I}$ such that $|B|<|A|$, then there exists an $x\in A\setminus B$ such that $B\cup\{x\}\in\mathcal{I}$.
+It turns out that given a set $S\subseteq E$, the maximal independent subsets of $S$ all have the same cardinality.
+This cardinality is called the rank of the set $S$, and is denoted by $r(S)$.
+The maximum size independent sets of $M$ are called bases, and the set of bases is denoted by $\mathcal{B}(M)$.
 The dual of a matroid $M$ is the matroid $M^{*}$ whose bases are $\{E\setminus B\ |\ B\in\mathcal{B}(M)\}$.
-For further details on matroids, see for instance [20]
-Given a graph $G=(V,E)$, the edge sets of the spanning forests of $G$ form the independent sets of a matroid $M_{G}$ called the cycle matroid of $G$.
-If $G$ is connected, then the basis of $M_{G}$ are the spanning trees of $G$.
+A circuit of a matroid is a minimal dependent set $A$, that is, each proper subset of $A$ is an independent set, but $A$ itself is not in $\mathcal{I}$.
+For further details on matroids, see for instance [20].
+This paper assumes very little familiarity with matroid theory, and we offer the following two families of examples for those unfamiliar with the basics.
+Let $U_{r,n}$ denote the uniform matroid of rank $r$ on $n$ elements: here the set $\mathcal{I}$ consists of the subsets of $\{1,2,\dots,n\}$ that have size at most $r$.
+Clearly, $\mathcal{I}$ satisfies all three axioms.
+Here the bases are the subsets of $\{1,2,\dots,n\}$ that have size exactly $r$.
+The circuits are the sets of size $r+1$, and the dual matroid is $U_{n-r,n}$.
+Another important family of matroids are the so-called cycle matroids.
+Given a graph $G=(V,E)$ we can consider the matroid $M_{G}$ on the ground set $E$, where $A\subseteq E$ is independent if it does not contain a cycle, that is, it is a forest.
+It is clear that it satisfies (i) and (ii), and a little inspection shows that it also satisfies (iii).
+This is called the cycle matroid of $G$, and is denoted by $M_{G}$.
+One can see that if $G$ is connected, then the bases of $M_{G}$ are the spanning trees of $G$.
+In general, the rank of a set $A\subseteq E$ is v(G)k(A)v(G)-k(A), where $k(A)$ is the number of connected components of the subgraph $(V,A)$.
+A circuit of $M_{G}$ corresponds to the edge set of a cycle of $G$.
+We believe that understanding uniform matroids and cycle matroids is sufficient to have a good grasp on the content of this paper.
 One can define the Tutte polynomial of a matroid as
@@
 Equivalently, a coloop is an element that is in every base of $M$.
-For a cycle matroid $M_{G}$, loops correspond to loop edges and coloops correspond to bridges in the graph $G$.
-Hence it was suggested that the inequalities
+For a cycle matroid $M_{G}$, loops correspond to loop edges, and coloops correspond to bridges in the graph $G$.
+Hence, it was suggested that the inequalities
 $\max(T_{M}(2,0),T_{M}(0,2)\geqslant T_{M}(1,1),$
@@
 may hold true for all matroids $M$ without loops and coloops.
-(These versions appear explicitly in [10], but were treated much earlier without explicitly calling them conjectures.) Note that for general matroids, all these versions are equivalent in the following sense: if one of them is true for all matroids, then the others are also true for all matroids.
+(These versions appear explicitly in [10], but were treated much earlier without explicitly calling them conjectures.) Note that for general matroids, all these variants are equivalent in the following sense: if one of them is true for all matroids, then the others are also true for all matroids.
 Applying the maximum version to $M\oplus M^{*}$ with $M^{*}$ being the dual of $M$ leads to the multiplicative version of the conjecture.
@@
 2.2. Permutation Tutte polynomial
-The proof of Theorem 1.3 heavily relies on the theory of permutation Tutte polynomial.
-The idea is that Tutte polynomial $T_{G}(x,y)$ can be written as a sum of permutation Tutte polynomials $\widetilde{T}_{H_{j}}(x,y)$ for certain bipartite graphs $H_{j}$.
-As a consequence certain inequalities valid for the permutation Tutte polynomial transfers to the Tutte polynomial.
+The proof of Theorem 1.3 heavily relies on the theory of the permutation Tutte polynomial.
+The idea is that the Tutte polynomial $T_{G}(x,y)$ can be written as a sum of the permutation Tutte polynomials $\widetilde{T}_{H_{j}}(x,y)$ for certain bipartite graphs $H_{j}$.
+As a consequence certain inequalities valid for the permutation Tutte polynomial transfer to the Tutte polynomial.
 Definition 2.1 ([1]).
@@
 $\pi(i)>\max_{j\in N_{H}(i)}\pi(j),$
-where the maximum over an empty set is set to be $\infty$- ∞. Similarly, we say that vertex $j\in B$ is externally active if
+where the maximum over an empty set is set to be -$\infty$. Similarly, we say that vertex $j\in B$ is externally active if
 $\pi(j)>\max_{i\in N_{H}(j)}\pi(i).$
 Let $\mathrm{ia}(\pi)$ and $\mathrm{ea}(\pi)$ be the number of internally and externally active vertices in $A$ and $B$, respectively. Let
-$\widetilde{T}_{H}(x,y)=\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}(\pi)}y^{\mathrm{ea}(\pi)}.$
-We will call $\widetilde{T}_{H}(x,y)$ the permutation Tutte polynomial of $H$.
+$\widetilde{T}_{H}(x,y)=\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}(\pi)}y^{\mathrm{ea}(\pi)},$
+where $S_{m}$ denotes the set of all permutations on $m$ elements. We will call $\widetilde{T}_{H}(x,y)$ the permutation Tutte polynomial of $H$.
 The above definition is motivated by the following theorem of Tutte.
@@
 Label the edges with $1,2,\dots,m$ arbitrarily.
-In the case of a spanning tree $T$ of $G$, let us call an edge $e\in E(T)$ internally active if $e$ has the largest label among the edges in the cut determined by $T$ and $e$ by removing $e$ from $T$.
-Let us call an edge $e\notin E(T)$ externally active if $e$ has the largest label among the edges in the cycle determined by $T$ and $e$ by adding $e$ to $T$.
+In the case of a spanning tree $T$ of $G$, let us call an edge $e\in E(T)$ internally active if $e$ has the largest label among the edges $f^{\prime}\in E(G)$ that are going between the connected components of the graph TeT-e.
+Let us call an edge $f\notin E(T)$ externally active if $e$ has the largest label among the edges in the cycle determined by $T$ and $f$ by adding $f$ to $T$.
 Let $\mathrm{ia}(T)$ and $\mathrm{ea}(T)$ be the number of internally and externally active edges, respectively.
@@
 $T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}(T)}y^{\mathrm{ea}(T)},$
-where the summation goes for all spanning trees of $G$.
+where the summation runs over all spanning trees of $G$.
 Theorem 2.2 was originally a definition for the Tutte polynomial [22].
+This theorem or definition naturally extends to matroids even though Tutte’s original paper only concerned graphs.
 This characterization of the Tutte polynomial immediately shows that the coefficients of the Tutte polynomial are non-negative.
-In this theorem, we are restricted to the same labelling of the edges for all spanning trees.
-For those who have never seen this definition before, it might be very surprising that the Tutte polynomial is independent of the actual choice of the labelling.
+In this theorem, we are restricted to the same labeling of the edges for all spanning trees.
+For those who have never seen this definition before, it might be very surprising that the Tutte polynomial is independent of the actual choice of the labeling.
 To explain the connection between $T_{G}(x,y)$ and $\widetilde{T}_{H}(x,y)$, we need the concept of the local basis exchange graph.
@@
 One bipartite class consists of the edges of $T$, the other consists of the edges of $E\setminus T$, and we connect a spanning tree edge $e$ with a non-edge $f$ if $f$ is in the cut determined by $e$ and $T$, equivalently, $e$ is in the cycle determined by $f$ and $T$.
-Clearly, this definition works for general matroids and their basis: if $A$ is a basis, then $e\in A$ and $f\in E\setminus A$ are adjacent in the bipartite graph $H[A]$ if $A-e+f$ is again a basis.
-Figure 1 depicts a graph $G$ with a spanning tree $T$ and the bipartite graph $H[T]$ obtained from $T$.
-For a fixed labelling of the edges of $G$, we get a labelling of the vertices of $H[T]$, and the internally (externally) active edges of $G$ correspond to internally (externally) active vertices of $H[T]$, so the two definitions of internal and external activity are compatible.
+This is also equivalent with Te+fT-e+f being a spanning tree of $G$ again.
+Clearly, this definition works for general matroids $M=(E,\mathcal{I})$ and their basis.
+If $A$ is a basis, then let $H[A]$ be the bipartite graph with $V(H[A])=E$, where one part consists of the elements of $A$, the other part consists of $E\setminus A$, and $e\in A$ and $f\in E\setminus A$ are adjacent in the bipartite graph $H[A]$ if Ae+fA-e+f is again a basis.
+Figure 1 shows a graph $G$ with a spanning tree $T$ and the bipartite graph $H[T]$ obtained from $T$.
+For a fixed labeling of the edges of $G$, we get a labeling of the vertices of $H[T]$, and the internally (externally) active edges of $G$ correspond to internally (externally) active vertices of $H[T]$, so the two definitions of internal and external activity are compatible.
 The following lemma is crucial for us, so we even included its proof.
-11112222333344446666555577778888
-11112222333355557777444466668888
+1122334466557788
+1122335577446688
 Figure 1. Example for a graph $G$ and the local basis exchange graph $H[T]$ obtained from a spanning tree $T$.
 Lemma 2.4 (Beke, Csáji, Csikvári, Pituk [1]).
-Let $G$ be a graph. For each spanning tree $T$ of $G$, let $H[T]$ be the local basis exchange graph with respect to $T$. Then
-$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y),$
-where the sum is over the set of spanning trees $\mathcal{T}(G)$ of $G$.
+Let $M$ be a matroid. For each basis $A$ of $M$, let $H[A]$ be the local basis exchange graph with respect to $A$. Then
+$T_{M}(x,y)=\sum_{A\in\mathcal{B}(M)}\widetilde{T}_{H[A]}(x,y),$
+where the sum runs over the set of bases $\mathcal{B}(M)$ of $M$.
 Proof.
-For a fixed spanning tree $T$ and a permutation $\pi$ of the edges, the internally and externally active edges correspond to the internally and externally active vertices of $H[T]$. Hence
-$T_{G}(x,y)=\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}.$
+For a fixed basis $A$ and a permutation $\pi$ of the edges, the internally and externally active edges correspond to the internally and externally active vertices of $H[A]$. Hence
+$T_{M}(x,y)=\sum_{A\in\mathcal{B}(M)}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}.$
 Now averaging it for all permutations $\pi\in S_{m}$ we get that
-$T_{G}(x,y)$ $=\frac{1}{m!}\sum_{\pi\in S_{m}}T_{G}(x,y)$
-$=\frac{1}{m!}\sum_{\pi\in S_{m}}\sum_{T\in\mathcal{T}(G)}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
-$=\sum_{T\in\mathcal{T}(G)}\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}_{H[T]}(\pi)}y^{\mathrm{ea}_{H[T]}(\pi)}$
-$=\sum_{T\in\mathcal{T}(G)}\widetilde{T}_{H[T]}(x,y).$
+$T_{M}(x,y)$ $=\frac{1}{m!}\sum_{\pi\in S_{m}}T_{M}(x,y)$
+$=\frac{1}{m!}\sum_{\pi\in S_{m}}\sum_{A\in\mathcal{B}(M)}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}$
+$=\sum_{A\in\mathcal{B}(M)}\frac{1}{m!}\sum_{\pi\in S_{m}}x^{\mathrm{ia}_{H[A]}(\pi)}y^{\mathrm{ea}_{H[A]}(\pi)}$
+$=\sum_{A\in\mathcal{B}(M)}\widetilde{T}_{H[A]}(x,y).$
 Remark 2.5.
-The local basis exchange graph $H[T]$ has an isolated vertex if and only if $G$ contains a bridge or a loop. Furthermore, $H[T]$ is connected if and only $G$ is $2$-connected.
-The following lemma enables us to study quadratic inequalities of the Tutte polynomial.
+The local basis exchange graph $H[A]$ has an isolated vertex if and only if $M$ contains a loop or a coloop.
+Furthermore, in case of a graph $G$ and a spanning tree $T$, the graph $H[T]$ is connected if and only if $G$ is $2$-connected.
+The following lemma is a variant of the transfer lemma [1] with the exact same proof. It enables us to study certain quadratic inequalities of the Tutte polynomial.
 Lemma 2.6 (Transfer lemma [1]).
-Let $x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}\geq 0$. Suppose that for any bipartite graph $H$, we have
-$\widetilde{T}_{H}(x_{1},y_{1})\widetilde{T}_{H}(x_{2},y_{2})\geq\widetilde{T}_{H}(x_{0},y_{0})^{2}.$
-Then for any graph $G$, we have
-$T_{G}(x_{1},y_{1})T_{G}(x_{2},y_{2})\geq T_{G}(x_{0},y_{0})^{2}.$
+Let $M$ be a matroid and $x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}\geqslant 0$. Suppose that for any basis $A$ of the matroid $M$, the local basis exchange graph $H[A]$ satisfies
+$\widetilde{T}_{H[A]}(x_{1},y_{1})\widetilde{T}_{H[A]}(x_{2},y_{2})\geqslant\widetilde{T}_{H[A]}(x_{0},y_{0})^{2},$
+then
+$T_{M}(x_{1},y_{1})T_{M}(x_{2},y_{2})\geqslant T_{M}(x_{0},y_{0})^{2}.$
 Remark 2.7.
-For us the following version of the transfer lemma will be a bit more convenient. If for any basis $A$ of a matroid $M$, the local basis exchange graph $H[A]$ satisfies that
-$\widetilde{T}_{H[A]}(x_{1},y_{1})\widetilde{T}_{H[A]}(x_{2},y_{2})\geq\widetilde{T}_{H[A]}(x_{0},y_{0})^{2},$
-then
-$T_{M}(x_{1},y_{1})T_{M}(x_{2},y_{2})\geq T_{M}(x_{0},y_{0})^{2}.$
-The proof of this claim is exactly the same as of the transfer lemma.
+In this paper, we aim to prove the inequality
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
+for various values of $x$. Observe that $\widetilde{T}_{H}(1,1)=1$, a direct consequence of the definition. So what we really need to prove is
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant 1.$
 A key example for bounding the permutation Tutte polynomial is the following theorem proved in [1].
 Theorem 2.8 ([1]).
-Let $H$ be an arbitrary bipartite graph, and let $d_{i}$ be the degree of vertex $i$. Suppose that $0\leq x\leq 1$ and $y\geq 1$ or $0\leq y\leq 1$ and $x\geq 1$. Then
-$\widetilde{T}_{H}(x,y)\geq\prod_{i\in A}\left(1+\frac{x-1}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(1+\frac{y-1}{d_{j}+1}\right).$
-The proof of Lemma 2.8 relies on following inequality of Harris that is also a special case of the FKG-inequality [11].
+Let $H=(A,B,E)$ be an arbitrary bipartite graph, and let $d_{i}$ be the degree of a vertex $i$.
+Suppose that $0\leqslant x\leqslant 1$ and $y\geqslant 1$, or $0\leqslant y\leqslant 1$ and $x\geqslant 1$.
+Then
+$\widetilde{T}_{H}(x,y)\geqslant\prod_{i\in A}\left(1+\frac{x-1}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(1+\frac{y-1}{d_{j}+1}\right).$
+The proof of Lemma 2.8 is based on the following inequality of Harris [12] that is also a special case of the FKG-inequality [11].
 Lemma 2.9 (Harris [12], Fortuin, Kasteleyn, Ginibre [11]).
-Suppose that $\mu=\mu_{1}\otimes\dots\otimes\mu_{N}$ is a product measure on a set $S_{1}\times S_{2}\times\dots\times S_{N}$, and $X_{1},\dots,X_{t}$ are non-negative monotone increasing functions in the sense that if $x_{i}\geq x_{i}^{\prime}$ for $i=1,\dots,N$, then for $1\leq j\leq t$ we have
-$X_{j}(x_{1},\dots,x_{N})\geq X_{j}(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
+Let $S_{1}\times S_{2}\times\dots\times S_{N}$ be a set, and suppose that $\mu=\mu_{1}\otimes\dots\otimes\mu_{N}$ is a product measure on it.
+Let $X_{1},\dots,X_{t}$ be non-negative monotone increasing functions in the sense that if $x_{i}\geqslant x_{i}^{\prime}$ for $i=1,\dots,N$, then for $1\leqslant j\leqslant t$ we have
+$X_{j}(x_{1},\dots,x_{N})\geqslant X_{j}(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
 Then
-$\mathbb{E}_{\mu}\left[\prod_{j=1}^{t}X_{j}\right]\geq\prod_{j=1}^{t}\mathbb{E}_{\mu}[X_{j}].$
-Furthermore, if $X$ is monotone increasing and $Y$ is monotone decreasing, then
-$\mathbb{E}[XY]\leq\mathbb{E}[X]\mathbb{E}[Y].$
-In this paper, we will use the Harris inequality for both $[0,1]^{N}$ and the set $\{0,1\}^{A}$ for some set $A$.
+$\mathbb{E}_{\mu}\left[\prod_{j=1}^{t}X_{j}\right]\geqslant\prod_{j=1}^{t}\mathbb{E}_{\mu}[X_{j}].$
+Moreover, let $Y$ be a monotone decreasing function, that is, for $x_{i}\geqslant x_{i}^{\prime}$ $(1\leqslant i\leqslant N)$ we have
+$Y(x_{1},\dots,x_{N})\leqslant Y(x_{1}^{\prime},\dots,x_{N}^{\prime}).$
+If $X$ is monotone increasing and $Y$ is monotone decreasing, then
+$\mathbb{E}_{\mu}[XY]\leqslant\mathbb{E}_{\mu}[X]\mathbb{E}_{\mu}[Y].$
+In this paper, we will use Harris’s inequality for both the product space of $[0,1]$ intervals, that is, $0,1]^{N}$, and for the discrete set $\{0,1\}^{A}$ for some set $A$.
 In what follows, we repeatedly use the following crucial idea to express $\widetilde{T}_{H}(x,y)$.
 We can create a random ordering of the vertices of $H$ as follows: for each vertex $i$ we choose a uniform random number $x_{i}$ from the interval $[0,1]$.
-The numbers $x_{i}$ then determine an ordering of the edges.
-The probability that two numbers are equal is 00.
+The numbers $x_{i}$ then determine the ordering of the vertices of $H$.
+The probability that two numbers are equal is 0.
 Lemma 2.10 ([1]).
@@
 $\mathbb{P}\left(I(A)=i,I(B)=j\right)=t_{i,j}(H).$
-In what follows we do a little trick.
-For $i\in A$ we generate $x_{i}\sim U(0,1)$ as before, but for $j\in B$ we actually first generate a uniformly random number $y_{j}$ from $[0,1]$ and let $x_{j}=1-y_{j}$.
-The role of this trick will be apparent soon.
 For $i\in A$, let us introduce the random variable
-$X_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$
+$Z_{i}(x_{i},\{x_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}>x_{i}.\end{array}\right$.
 and for $j\in B$, let
-$Y_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
+$Z_{j}(\{x_{i}\}_{i\in A},x_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant x_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>x_{j}.\end{array}\right$.
+As a consequence of Lemma 2.10 we have
+(2.1) $\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
+The problem with the random variables $Z_{v}(\underline{x})$ is that they are neither monotone increasing, nor monotone decreasing.
+In what follows, we do a little trick.
+For $i\in A$ we generate $x_{i}\sim U(0,1)$ as before, but for $j\in B$ we first generate a uniform random number $y_{j}$ from $[0,1]$ and let $x_{j}=1-y_{j}$.
+The role of this trick will become apparent soon.
+For $i\in A$, let us introduce the random variable
+$X_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right$.
+and for $j\in B$, let
+$Y_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>1-y_{j}.\end{array}\right$.
 Lemma 2.11.
@@
 $\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{i\in A}X_{i}\cdot\prod_{j\in B}Y_{j}\right].$
-(b1) If $x\geq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function for each $i\in A$.
-(b2) If $0\leq x\leq 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function for each $i\in A$.
-(b3) For $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
-(b4) Finally, for $0\leq y\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing for each $j\in B$.
-Remark 2.12.
-This lemma is the reason why we generated $x_{j}$ by $x_{j}=1-y_{j}$ for $j\in B$. If we consider the function
-$X^{\prime}_{i}(x_{i},\{x_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}x_{j}>x_{i}.\end{array}\right.$
-instead of $X_{i}(x_{i},\{y_{j}\}_{j\in B})$, this would be neither increasing, nor decreasing.
-We will also need a very little extension of the above ideas, where each vertex gets its own activity: let For $i\in A$, let us introduce the random variable
-$\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x^{(i)}&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leq x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right.$
+(b1) If $x\geqslant 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function for each $i\in A$.
+(b2) If $0\leqslant x\leqslant 1$, then $X_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function for each $i\in A$.
+(b3) For $0\leqslant y\leqslant 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ increases monotonically for each $j\in B$.
+(b4) Finally, for $y\geqslant 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ decreases monotonically for each $j\in B$.
+We will also need a slight extension of the above ideas, where each vertex gets its own activity. For $i\in A$, let us introduce the random variable
+$\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})=\left\{\begin{array}[]{ll}x^{(i)}&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})\leqslant x_{i},\\ 1&\mbox{if}\ \max_{j\in N_{H}(i)}(1-y_{j})>x_{i}.\end{array}\right$.
 and for $j\in B$, let
-$\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y^{(j)}&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leq 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\geq 1-y_{i}.\end{array}\right.$
-The following lemma is just a trivial extension of the previous lemma together with Harris’ inequality.
+$\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})=\left\{\begin{array}[]{ll}y^{(j)}&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}\leqslant 1-y_{j},\\ 1&\mbox{if}\ \max_{i\in N_{H}(j)}x_{i}>1-y_{j}.\end{array}\right$.
+The following lemma is just a trivial extension of the previous lemma together with Harris’s inequality.
 Lemma 2.12.
-(a1) If $i\in A$ and $x^{(i)}\geq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function.
-(a2) If $i\in A$ and $0\leq x^{(i)}\leq 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function.
-(a3) For $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing for each $j\in B$.
-(a4) Finally, if $j\in B$ and $0\leq y^{(j)}\leq 1$ the function $Y_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing.
-(b) If $x^{(i)}\geq 1$ for all $i\in A$ and $0\leq y^{(j)}\leq 1$, then
-$\mathbb{E}\left[\prod_{i\in A}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]\geq\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]=\prod_{i\in A}\left(\frac{x^{(i)}}{d_{i}+1}+\frac{d_{i}}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(\frac{y^{(j)}}{d_{j}+1}+\frac{d_{j}}{d_{j}+1}\right).$
+(a1) If $i\in A$ and $x^{(i)}\geqslant 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone increasing function.
+(a2) If $i\in A$ and $0\leqslant x^{(i)}\leqslant 1$, then $\widehat{X}_{i}(x_{i},\{y_{j}\}_{j\in B})$ is a monotone decreasing function.
+(a3) If $j\in B$ and $0\leqslant y^{(j)}\leqslant 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone increasing.
+(a4) Finally, if $j\in B$ and $y^{(j)}\geqslant 1$ the function $\widehat{Y}_{j}(\{x_{i}\}_{i\in A},y_{j})$ is monotone decreasing.
+(b) If $x^{(i)}\geqslant 1$ for all $i\in A$ and $0\leqslant y^{(j)}\leqslant 1$ for all $j\in B$, then
+$\mathbb{E}\left[\prod_{i\in A}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]\geqslant\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]=\prod_{i\in A}\left(\frac{x^{(i)}}{d_{i}+1}+\frac{d_{i}}{d_{i}+1}\right)\cdot\prod_{j\in B}\left(\frac{y^{(j)}}{d_{j}+1}+\frac{d_{j}}{d_{j}+1}\right).$
 We will use one more lemma from the paper [1], namely the gluing lemma. Originally, this lemma used the condition that $H_{1}$ and $H_{2}$ are trees, but the proof never used this condition.
 Lemma 2.13 (Gluing lemma [1]).
-Let $x\geq 1$ and $0\leq y\leq 1$.
-Let $H_{1}$ be a rooted graph with root vertex $v_{1}$.
-Let $H_{2}$ be another rooted graph with root vertex $v_{2}$.
+Let $x\geqslant 1$ and $0\leqslant y\leqslant 1$.
+Let $H_{1}$ and $H_{2}$ be rooted bipartite graphs with root vertices $v_{1}$ and $v_{2}$, respectively.
 Let $H$ be obtained from $H_{1}$ and $H_{2}$ by identifying $v_{1}$ and $v_{2}$ in the union of $H_{1}$ and $H_{2}$.
@@
 (a) If $v\in A$, then
-$x\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
+$x\widetilde{T}_{H}(x,y)\geqslant\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
 (b) If $v\in B$, then
-$\widetilde{T}_{H}(x,y)\geq\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
+$\widetilde{T}_{H}(x,y)\geqslant\widetilde{T}_{H_{1}}(x,y)\widetilde{T}_{H_{2}}(x,y).$
 (c) In particular,
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\frac{1}{x}(\widetilde{T}_{H_{1}}(x,0)\widetilde{T}_{H_{1}}(0,x))(\widetilde{T}_{H_{2}}(x,0)\widetilde{T}_{H_{2}}(0,x))$
-3. Warm-up: asymptotic computation of some permutation Tutte polynomial
-In this section we motivate an important technique of this paper by computing the asymptotic value of the permutation Tutte polynomial of certain bipartite graphs.
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\frac{1}{x}(\widetilde{T}_{H_{1}}(x,0)\widetilde{T}_{H_{1}}(0,x))(\widetilde{T}_{H_{2}}(x,0)\widetilde{T}_{H_{2}}(0,x)).$
+3. Warm-up: asymptotic computation of some permutation Tutte polynomials
+In this section, we motivate an important technique of this paper by computing the asymptotic values of the permutation Tutte polynomial of certain bipartite graphs.
 Let us start with a very simple example, the complete bipartite graph.
@@
 We have
-[…]
+$\widetilde{T}_{K_{a,b}}(x,0)=ab\int_{0}^{1}\int_{s}^{1}s^{b-1}(s+x(t-s))^{a-1}\,dt\,ds.$
 Proof.
-We generate the random permutation on $V(H)=A\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s.
+Let $K_{a,b}=(A,B,E)$ with $|A|=a,|B|=b$.
+As before we generate the random permutation on $A\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V(K_{a,b})$, and then we take the relative order of $x_{v}$’s.
 Let $t=\max_{v\in A}x_{v}$ and $s=\max_{v\in B}x_{v}$.
-We have $ab$ choices for the vertices that take the value $t$ and $s$, let these vertices be $v_{A}$ and $v_{B}$.
-If $s>t$, then the permutation has weight 00 as $v_{B}$ is an active vertex.
+Recall that from equation 2.1 we have
+$\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
+We will refer to $\prod_{v\in V(H)}Z_{v}(\underline{x})$ as the weight corresponding to the permutation determined by $\underline{x}$ as it only depends on the ordering of $(x_{v})_{v\in V}$.
+Our plan is to compress this integral using $s$ and $t$.
+We have $ab$ choices for the vertices that take the values $t$ and $s$, let these vertices be $v_{A}$ and $v_{B}$.
+If $s>t$, then the permutation has weight 0 as $v_{B}$ is an active vertex.
 If $t>s$, then no vertex can be active in $B$.
-For each $w\in A$ that is not equal to $v_{A}$ two things can happen: if $x_{w}<s$, then $w$ is not active, so $X_{w}=1$, or $s<x_{w}<t$ and then $w$ is active so $X_{w}=x$.
-For a vertex $w\in B$ we simply need to have $x_{w}<s$ and $w$ will not be active.
+For each $w\in A$ that is not equal to $v_{A}$ two things can happen: if $x_{w}<s$, then $w$ is not active, so $Z_{w}(\underline{x})=1$, or $s\leqslant x_{w}\leqslant t$ and then $w$ is active so $Z_{w}(\underline{x})=x$.
+This shows that as $x_{w}$ runs over the interval $[0,t]$ its contribution to the integral is s+x(ts)s+x(t-s).
+For a vertex $w\in B$ we simply need to have $x_{w}<s$ and $w$ will not be active, so we integrate $1$ on the whole interval $[0,s]$.
 The integral formula then follows.
@@
 Clearly, at the maximum we have $t=1$ and we simply need to maximize $s^{\beta}(s+x(1-s))^{\alpha}$.
-This turns out to be at $s=\min\left(1,\frac{\beta x}{\beta-1}\right)$.
+This turns out to be at $s=\min\left(1,\frac{\beta x}{x-1}\right)$.
 If $\beta<\frac{x-1}{x}$, then the growth constant is
-$\left(\frac{\beta x}{\beta-1}\right)^{\beta}\left(x+(1-x)\frac{\beta x}{x-1}\right)^{\alpha}=\alpha^{\alpha}(1-\alpha)^{1-\alpha}\frac{x}{(x-1)^{1-\alpha}}.$
-If $\beta\geq\frac{x-1}{x}$, then the exponential growth constant is simply $1$.
+$\left(\frac{\beta x}{x-1}\right)^{\beta}\left(x+(1-x)\frac{\beta x}{x-1}\right)^{\alpha}=\alpha^{\alpha}(1-\alpha)^{1-\alpha}\frac{x}{(x-1)^{1-\alpha}}.$
+If $\beta\geqslant\frac{x-1}{x}$, then the exponential growth constant is simply $1$.
 Let us consider the graph $H_{a,b,c}$ introduced in [1]: we start with a complete bipartite graph $K_{a,b}$ with vertex set $A\cup B$, and then attach $c$ pendant leaves to $c$ distinct vertices of $B$, let $C$ be the set of these leaf vertices.
@@
 It turns out that for even $n$ there are matroids for which all local basis exchange graphs are isomorphic to $H_{n,n,n}$.
-Indeed, all local basis exchange graphs of the matroid $U^{(2)}_{\frac{3}{2}n,n}$ are isomorphic to $H_{n,n,n}$, where $U_{\frac{3}{2}n,n}$ is the uniform matroid of rank $n$ on $\frac{3}{2}n$ elements, and the matroid $M^{(2)}$ is obtained from $M$ by replacing each element of $M$ with two parallel elements.
-Furthermore, we have
-$\widetilde{T}_{H_{n,n,n}}(2,0)\widetilde{T}_{H_{n,n,n}}(2,0)<\widetilde{T}_{H_{n,n,n}}(1,1)^{2}$
+Indeed, let us first consider the uniform matroid $U_{n,\frac{3}{2}n}$ of rank $n$ on $\frac{3}{2}n$ elements.
+Then let us replace all elements of $U_{n,\frac{3}{2}n}$ with $2$ parallel elements.
+Let us denote the resulting matroid by $U^{(2)}_{n,\frac{3}{2}n}$.
+(In general, if we replace each element of a matroid $M$ with $k$ parallel elements, then we denote by $M^{(k)}$ the obtained matroid.) The resulting matroid has $3n$ elements and rank $n$.
+If $B$ is a basis of the matroid $U^{(2)}_{n,\frac{3}{2}n}$, then $C$ will be the set of corresponding parallel elements, and $A$ will be the remaining elements.
+So the local basis exchange graph will be the dual bipartite graph $H^{*}_{n,n,n}=(B,A\cup C,E)$.
+For the dual matroid of $U^{(2)}_{n,\frac{3}{2}n}$ the local basis exchange graph will be exactly $H_{n,n,n}=(A\cup C,B,E)$.
+Below we sketch an alternative proof for the fact that
+$\widetilde{T}_{H_{n,n,n}}(2,0)\widetilde{T}_{H_{n,n,n}}(0,2)<\widetilde{T}_{H_{n,n,n}}(1,1)^{2}$
 for large enough $n$.
-We can now give a simple proof of this fact.
 Theorem 3.3.
 We have
-$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}=\max_{s\in[0,1]}\left(s+x(1-s)\right)\left(xs+(1-x)\frac{s^{2}}{2}\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}&\text{if}\ \ x\geq\sqrt{3},\\ \frac{1}{2}(x+1)&\text{if}\ \ 1<x\leq\sqrt{3}\end{cases}$
+$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}=\max_{s\in[0,1]}\left(s+x(1-s)\right)\left(xs+(1-x)\frac{s^{2}}{2}\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}&\text{if}\ \ x\geqslant\sqrt{3},\\ \frac{1}{2}(x+1)&\text{if}\ \ 1<x\leqslant\sqrt{3},\end{cases}$
 and
-$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}=\max_{s\in[0,1]}t\left(\frac{t^{2}}{2}+\left(\frac{1}{2}-\frac{t^{2}}{2}\right)x\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}&\text{if}\ \ x\geq\frac{3}{2},\\ \frac{1}{2}&\text{if}\ \ 1<x\leq\frac{3}{2}.\end{cases}$
-In particular, if $x\geq\sqrt{3}$ we have
+$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}=\max_{t\in[0,1]}t\left(\frac{t^{2}}{2}+\left(\frac{1}{2}-\frac{t^{2}}{2}\right)x\right)=\begin{cases}\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}&\text{if}\ \ x\geqslant\frac{3}{2},\\ \frac{1}{2}&\text{if}\ \ 1<x\leqslant\frac{3}{2}.\end{cases}$
+In particular, if $x\geqslant\sqrt{3}$ we have
 $\lim_{n\to\infty}\left(\widetilde{T}_{H_{n,n,n}}(x,0)\widetilde{T}_{H_{n,n,n}}(0,x)\right)^{1/n}=\left(\frac{x^{3}}{9(x-1)}\right)^{3/2}.$
 Sketch of the proof..
-As before we generate the random permutation on $V(H)$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let
+As before we generate the random permutation on
+$V(H)=(A\cup C)\cup B$ by first generating an $x_{v}\in(0,1)$ uniformly at random for all $v\in V$, and then we take the relative order of $x_{v}$’s. Let
 $t=\max_{v\in A}x_{v}\ \ \ \text{and}\ \ \ s=\max_{v\in B}x_{v}.$
-Let $v_{A}$ and $v_{B}$ be the vertices, where these maximum are achieved.
-(Note that $t$ is only maximum on $A$ and not on $A\cup C$.) If $s<t$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
-[…]
-Here $s+x(t-s)$ is the contribution of a vertex $w\in A$ that is not $v_{A}$.
+Let $v_{A}$ and $v_{B}$ be the vertices, where these maximums are achieved. (Note that $t$ is only maximum on $A$ and not on $A\cup C$.) As before we use equation 2.1:
+$\widetilde{T}_{H}(x,y)=\mathbb{E}\left[\prod_{v\in V(H)}Z_{v}(\underline{x})\right]=\int_{[0,1]^{V(H)}}.\prod_{v\in V(H)}Z_{v}(\underline{x})\ d\underline{x}.$
+and we aim to compress this integral using $s$ and $t$.
+If $s<t$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
+(3.1) $I_{1}:=n^{2}\int_{0}^{1}\int_{s}^{1}(s+x(t-s))^{n-1}\left(\int_{0}^{s}(s_{i}+x(1-s_{i}))\,ds_{i}\right)^{n-1}(s+x(1-s))\,dtds.$
+Here s+x(ts)s+x(t-s) is the contribution of a vertex $w\in A$ that is not $v_{A}$.
 If $x_{w}=s_{i}<s$ for some $w\in B$ that is not $v_{B}$, then the attached leaf in $C$ has contribution $s_{i}+x(1-s_{i})$.
-(Note that attached leaf vertex $u$ can have a value bigger than $t$ as $t$ was only a maximum among the vertices of $A$.) The last term $s+x(1-s)$ is simply the contribution of the leaf attached to $v_{B}$.
+(Note that attached leaf vertex $u$ can have a value bigger than $t$ as $t$ was only a maximum among the vertices of $A$.) The last term s+x(1s)s+x(1-s) is simply the contribution of the leaf attached to $v_{B}$.
 If $t<s$, then the contribution of this case to $\widetilde{T}_{H_{n,n,n}}(x,0)$ is
-[…]
-Here the contribution of a vertex $w\in A$ is simply $t$ as it can never be active since $s<t$.
+(3.2) $I_{2}:=n^{2}\int_{0}^{1}\int_{s}^{1}t^{n-1}\left(\int_{0}^{t}(s_{i}+x(1-s_{i}))\,ds_{i}+\int_{t}^{s}x(1-s_{i})\,ds_{i}\right)^{n-1}x(1-s)\,dtds.$
+Note that we can assume that none of the vertices of $B$ are active since the contribution of such a term would be 0.
+Here, the contribution of a vertex $w\in A\setminus\{v_{A}\}$ is simply $t$ as it can never be active since $t<s$.
 If we have a vertex $w\in B$ that is not $v_{B}$ with $x_{w}=s_{i}<s$, then we need to distinguish two cases.
 If $s_{i}<t$, then $w$ is not active, and so the attached leaf has a contribution $s_{i}+x(1-s_{i})$.
-If $s_{i}>t$ then, in order to make $w$ inactive the attached leaf $u$ should have a value $x_{u}>s_{i}$, but then it is automatically active.
-The contribution of the leaf attached to $v_{B}$ is $x(1-s)$ as it has to be active to make $v_{B}$ inactive.
+If $s_{i}>t$ then, in order to make $w$ inactive the attached leaf $u$ should have a value $x_{u}>s_{i}$, but then it is automatically active and its contribution is $x(1-s_{i})$.
+The contribution of the leaf attached to $v_{B}$ is x(1s)x(1-s) as it has to be active to make $v_{B}$ inactive.
 We have
-$\lim_{n\to\infty}I_{1}^{1/n}$ $=\max_{0\leq s\leq t\leq 1}(s+x(t-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
-$=\max_{0\leq s\leq 1}(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
-The function $(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$ has either maximum at $\left(1-\frac{1}{\sqrt{3}}\right)\frac{x}{x-1}$ if this is less than $1$, or at $s=1$.
+$\lim_{n\to\infty}I_{1}^{1/n}$ $=\max_{0\leqslant s\leqslant t\leqslant 1}(s+x(t-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
+$=\max_{0\leqslant s\leqslant 1}(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$
+The function $(s+x(1-s))\left(xs+(1-x)\frac{s^{2}}{2}\right)$ has either a maximum at $\left(1-\frac{1}{\sqrt{3}}\right)\frac{x}{x-1}$ if this is less than $1$, or at $s=1$.
 In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3}}{x-1}$.
@@
 On the other hand,
-$\lim_{n\to\infty}I_{2}^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left((1-x)\frac{t^{2}}{2}+xt+x(s-t)-x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
-$=\max_{0\leq t\leq s\leq 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
-$=\max_{0\leq s\leq 1}\left(x\left(s-\frac{s^{2}}{2}\right)+\frac{1}{2}\right)$
-$=\frac{1}{2}(x+1),$
-so this is always at most $\lim_{n\to\infty}I_{1}^{1/n}$.
-Then we have
+$\lim_{n\to\infty}I_{2}^{1/n}=\max_{0\leqslant t\leqslant s\leqslant 1}t\left((1-x)\frac{t^{2}}{2}+xt+x(s-t)-x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
+Observe that at $s=t=1$ its value is $\frac{1}{2}(x+1)$. On the other hand, this is the maximum as dropping the condition $t\leqslant s$ would lead to the following chain of inequalities
+$\lim_{n\to\infty}I_{2}^{1/n}$ $=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
+$\leqslant\max_{0\leqslant t,s\leqslant 1}t\left(xs-x\frac{s^{2}}{2}+\frac{t^{2}}{2}\right)$
+$=\max_{0\leqslant s\leqslant 1}\left(x\left(s-\frac{s^{2}}{2}\right)+\frac{1}{2}\right)$
+$=\frac{1}{2}(x+1).$
+So $\lim_{n\to\infty}I_{2}^{1/n}=\frac{1}{2}(x+1)$ which is always at most $\lim_{n\to\infty}I_{1}^{1/n}$. Then we have
 $\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(x,0)^{1/n}$ $=\lim_{n\to\infty}(I_{1}+I_{2})^{1/n}$
@@
 $=\lim_{n\to\infty}I_{1}^{1/n}.$
-This proves the first statement. The second statement is actually simpler. We have
-[…]
-[…]
+This proves the first statement.
+The statement concerning $\widetilde{T}_{H_{n,n,n}}(0,x)$ is actually simpler. As before, let
+$t=\max_{v\in A}x_{v}\ \ \ \text{and}\ \ \ s=\max_{v\in B}x_{v},$
+and let $v_{A}$ and $v_{B}$ be the vertices, where these maximums are achieved.
+Observe that we only have to consider the case $t\leqslant s$, otherwise the vertex $v_{A}$ is active and $X_{v_{A}}=0$.
+If $t\leqslant s$, then for any $v\in A$ different from $v_{A}$ the value $x_{v}$ can be anything between $[0,t]$ without $v$ being active.
+If for a vertex $v\in B$ different from $v_{B}$ we have $x_{v}=s_{i}$, then its unique neighbor in $C$ cannot take bigger value than $s_{i}$, otherwise it becomes active and the contribution of this permutation would be 0.
+If $s_{i}<t$, then $v$ is not active, while if $s_{i}\in(s,t)$, then $v$ is active.
+Putting all these together, we get that
+$\widetilde{T}_{H_{n,n,n}}(0,x)$ $=n^{2}\int_{0}^{1}\int_{0}^{s}t^{n-1}\left(\int_{0}^{t}s_{i}\,ds_{i}+x\int_{t}^{s}s_{i}\,ds_{i}\right)^{n-1}s\,dtds$
+$=n^{2}\int_{0}^{1}\int_{0}^{s}t^{n-1}\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)^{n-1}s\,dtds.$
 From this we get that
-$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}$ $=\max_{0\leq t\leq s\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
-$=\max_{0\leq t\leq 1}t\left(\frac{t^{2}}{2}+x\left(\frac{1}{2}-\frac{t^{2}}{2}\right)\right)$
-This expression has either maximum at $x=\left(\frac{x}{3(x-1)}\right)^{1/2}$ if $x\geq\frac{3}{2}$ or at $1$ if $x\leq\frac{3}{2}$.
-In the first case its value is $\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}$.
+$\lim_{n\to\infty}\widetilde{T}_{H_{n,n,n}}(0,x)^{1/n}$ $=\max_{0\leqslant t\leqslant s\leqslant 1}t\left(\frac{t^{2}}{2}+x\left(\frac{s^{2}}{2}-\frac{t^{2}}{2}\right)\right)$
+$=\max_{0\leqslant t\leqslant 1}t\left(\frac{t^{2}}{2}+x\left(\frac{1}{2}-\frac{t^{2}}{2}\right)\right)$
+This expression has a maximum either at $t=\left(\frac{x}{3(x-1)}\right)^{1/2}$ if $x\geqslant\frac{3}{2}$ or at $1$ if $x\leqslant\frac{3}{2}$.
+In the first case, its value is $\frac{1}{3\sqrt{3}}\frac{x^{3/2}}{(x-1)^{1/2}}$.
 In the second case, its value is $\frac{1}{2}$.
-This completes the sketch of the proof.
+Since $\widetilde{T}_{H}(1,1)=1$ for any bipartite graph $H$ by definition, this completes the sketch of the proof.
 ∎
 Remark 3.4.
-It is clear from the above examples that even though the permutation Tutte polynomial might be complicated still it is possible to compute the exponential growth of it for many families of bipartite graphs.
-Another important take-away that it is often possible to understand those permutations that contributes the most weight to the permutation Tutte polynomial, and concentrating for these permutations might give a better lower or upper bound for the permutation Tutte polynomial.
+There are two takeaways from this section.
+First, sometimes it is possible to compute the exponential growth of the permutation Tutte polynomial without explicitly computing it.
+Second, it is often possible to understand those permutations that contribute the most weight to the permutation Tutte polynomial.
+Concentrating on these permutations might give a better lower or upper bound on the permutation Tutte polynomial.
 4. The main lemma
-In this section we give the main ingredient of the proofs of Theorems 1.3 and 1.4.
+In this section, we give the main ingredient of the proofs of Theorems 1.3 and 1.4.
 Theorem 4.1.
-Let $H=(A,B,E)$ be a bipartite graph. For $x\geq 2$ and a fixed $0\leq s\leq 1$ let us introduce the function
+Let $H=(A,B,E)$ be a bipartite graph. For $x\geqslant 2$ and a fixed $0\leqslant s\leqslant 1$ let us introduce the function
 $\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}.$
 Then
-$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
+$\widetilde{T}_{H}(x,0)\geqslant\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
 Proof.
@@
 $\widetilde{T}_{H}(x,0)=\int_{[0,1]^{A}}\int_{[0,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
-The first idea is that we only take those values for which $x_{u}=1-y_{u}\leq s$ for $u\in B$. Hence
-$\widetilde{T}_{H}(x,0)\geq\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
-For each $v\in A$ we decompose the integral over $[0,1]$ to $[0,s]$ and $[s,1]$, this way we decomposed the integral $[0,1]^{A}$ to the sum of $2^{|A|}$ integrals.
-Since $x_{u}\leq s$ for $u\in B$ we immediately get that if $x_{v}\geq s$, then $X_{v}(\underline{x},\underline{y})=x$.
-Let $T\subseteq A$ be the set of those vertices $v\in A$ for which $x_{v}\geq s$.
+The first idea is that we only consider those $((x_{v})_{v\in A},(y_{u})_{u\in B})$ for which $x_{u}=1-y_{u}\leqslant s$ for $u\in B$. Hence
+$\widetilde{T}_{H}(x,0)\geqslant\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
+For each $v\in A$ we decompose the integral over $[0,1]$ to $[0,s]$ and $[s,1]$, this way we decomposed the integral $0,1]^{A}$ to the sum of $2^{|A|}$ integrals.
+Since $x_{u}\leqslant s$ for $u\in B$ we immediately get that if $x_{v}\geqslant s$, then $v$ is active and so $X_{v}(\underline{x},\underline{y})=x$.
+Let $T\subseteq A$ be the set of those vertices $v\in A$ for which $x_{v}\geqslant s$.
 Then
 $\int_{[0,1]^{A}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
-$=\sum_{T\subseteq A}(x(1-s))^{|T|}\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
+$=\sum_{T\subseteq A}(x(1-s))^{|T|}\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A\setminus T}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}.$
 Let $N(T)\subseteq B$ be the neighbors of $T$.
-If $u\in N(T)$, then $\widehat{Y}_{u}(\underline{x},\underline{y}):=Y_{u}(\underline{x},\underline{y})=1$, because the vertex $u$ cannot be active anymore.
+If $u\in N(T)$, then the vertex $u$ cannot be active anymore, and so let us define $\widehat{Y}_{u}(\underline{x},\underline{y}):=Y_{u}(\underline{x},\underline{y})=1$.
 Hence
-$I_{T}:=\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
-corresponds to $s^{|A\setminus T|+|B|}$ times a permutation Tutte polynomial, where an active vertex $i\in A\setminus T$ gets a weight $x^{(i)}=x$ while in $B$ an active vertex $j$ gets a weight $y^{(j)}=1$ if it is $N(T)$, and $y^{(j)}=0$ if it is $B\setminus N(T)$, so for this term we have a lower bound from Lemma 2.13:
+$I_{T}:=\int_{[0,s]^{A\setminus T}}\int_{[1-s,1]^{B}}\prod_{v\in A\setminus T}X_{v}(\underline{x},\underline{y})\cdot\prod_{u\in B}Y_{u}(\underline{x},\underline{y})d\underline{x}d\underline{y}$
+corresponds to $s^{|A\setminus T|+|B|}$ times a permutation Tutte polynomial on $(A\setminus T,B)$, where an active vertex $i\in A\setminus T$ gets a weight $x^{(i)}=x$ while in $B$ an active vertex $j$ gets a weight $y^{(j)}=1$ if it is $N(T)$, and $y^{(j)}=0$ if it is $B\setminus N(T)$, so for this term we have a lower bound from Lemma 2.12:
 $I_{T}$ $=s^{|A\setminus T|+|B|}\mathbb{E}\left[\prod_{i\in A\setminus T}\widehat{X}_{i}\cdot\prod_{j\in B}\widehat{Y}_{j}\right]$
-$\geq s^{|A\setminus T|+|B|}\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]$
+$\geqslant s^{|A\setminus T|+|B|}\prod_{i\in A}\mathbb{E}[\widehat{X}_{i}]\cdot\prod_{j\in B}\mathbb{E}[\widehat{Y}_{j}]$
 $=s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
 So far we got that
-$\widetilde{T}_{H}(x,0)$ $=\sum_{T\subseteq A}(x(1-s))^{|T|}I_{T}$
-$\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
+$\widetilde{T}_{H}(x,0)$ $\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}I_{T}$
+$\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot s^{|A\setminus T|+|B|}\prod_{v\in A\setminus T}\left(1+\frac{x-1}{d_{v}+1}\right)\prod_{u\in B\setminus N(T)}\left(1-\frac{1}{d_{u}+1}\right).$
 (4.1) $=\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
-Next let
+Next, let
 $Z=\prod_{v\in A}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)\right),$
@@
 $\mu(T)=\frac{1}{Z}(x(1-s))^{|T|}\prod_{A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right).$
-Note that $\mu$ is a product measure $\mu=\bigoplus_{v\in A}\mu_{v}$, where
+Note that $\mu$ is a product measure $\mu=\bigotimes_{v\in A}\mu_{v}$, where
 $\mu_{v}(0)=\frac{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s}{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)}\ \ \ \text{and}\ \ \ \mu_{v}(1)=\frac{x(1-s)}{\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s+x(1-s)}.$
-For each $u\in B$ consider the function
+For each $u\in B$ consider the function $f_{u}:\{0,1\}^{A}\to\mathbb{R}$ defined as
 $f_{u}(T)=\begin{cases}s&\text{if }u\in N(T),\\ \frac{d_{u}}{d_{u}+1}s&\text{if }u\not\in N(T).\end{cases}$
 Then we can rewrite inequality (4.1) as follows:
-$\widetilde{T}_{H}(x,0)$ $\geq\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
+$\widetilde{T}_{H}(x,0)$ $\geqslant\sum_{T\subseteq A}(x(1-s))^{|T|}\cdot\prod_{v\in A\setminus T}\left(\frac{d_{v}}{d_{v}+1}s+\frac{x}{d_{v}+1}s\right)\cdot\prod_{u\in B\setminus N(T)}\frac{d_{u}s}{d_{u}+1}\cdot\prod_{u\in N(T)}s$
 $=Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T).$
-All functions $f_{u}$ are monotone increasing on the space $\{0,1\}^{A}$ identified with the subsets $T$ of $A$, since adding vertices to $T$ can only increase $N(T)$, so by Harris’ inequality we have
-$Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T)\geq Z\prod_{u\in B}\left(\sum_{T\subseteq A}\mu(T)f_{u}(T)\right).$
+All functions $f_{u}$ are monotone increasing on the space $\{0,1\}^{A}$ identified with the subsets $T$ of $A$, since adding vertices to $T$ can only increase $N(T)$, so by Harris’s inequality (Lemma 2.9) we have
+$Z\sum_{T\subseteq A}\mu(T)\prod_{u\in B}f_{u}(T)\geqslant Z\prod_{u\in B}\left(\sum_{T\subseteq A}\mu(T)f_{u}(T)\right).$
 Here
@@
 Hence
-$\widetilde{T}_{H}(x,0)\geq\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
+$\widetilde{T}_{H}(x,0)\geqslant\prod_{v\in A}\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\cdot\prod_{u\in B}\left(s-\frac{s}{d_{u}+1}\prod_{v\in N(u)}\gamma_{x,s}(d_{v})\right).$
 By applying Theorem 4.1 for both $\widetilde{T}_{H}(x,0)$ and $\widetilde{T}_{H}(0,x)$ we get the following corollary.
@@
 We have
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right].$
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right].$
+Definition 4.3.
+Let us introduce the notation
+$\mathbb{G}\left(d,x,s,\{\gamma_{j}\}_{1\leqslant j\leqslant d}\right):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\prod_{j=1}^{d}\gamma_{j}\right).$
+Furthermore, let
+$G(d,x,s,\gamma):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right).$
+For any $0\leqslant\gamma<1$ we have
+$G(\infty,x,s):=\lim_{d\to\infty}G(d,x,s,\gamma)=(s+x(1-s))s.$
+Clearly, with this notation we can rewrite Corollary 4.2 as
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right).$
+Next, we prove some monotonicity properties of the function $\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$.
 Lemma 4.4.
-If $x\geq 1$ and $0\leq s\leq 1$, then the function $\gamma_{x,s}(d)$ is monotone decreasing in $d$. Furthermore, $\gamma_{x,s}(d)\leq s$.
+If $x\geqslant 1$ and $0\leqslant s\leqslant 1$, then the function $\gamma_{x,s}(d)$ is monotone decreasing in $d$. Furthermore, $\gamma_{x,s}(d)\leqslant s$.
 Proof.
@@
 In general, the derivative of a function $f(t)=\frac{at+b}{ct+d}$ is $\frac{ad-bc}{(ct+d)^{2}}$, so it is enough to check that
-$sx-sx(x(1-s)+s)=sx(1-x)(1-s)\leq 0$
-which is clearly true. The second statement follows from $d+x\leq(d+1)x$ as follows:
-$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}\leq\frac{(d+x)s}{(d+x)s+(d+x)(1-s)}=s.$
-Definition 4.4.
-Let
-$G(d,x,s,\gamma):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right).$
-For any $10\leq\gamma<$ 1 we have
-$G(\infty,x,s):=\lim_{d\to\infty}G(d,x,s,\gamma)=(s+x(1-s))s.$
+$sx-sx(x(1-s)+s)=sx(1-x)(1-s)\leqslant 0$
+which is clearly true. The second statement follows from $d+x\leqslant(d+1)x$ as follows:
+$\gamma_{x,s}(d)=\frac{(d+x)s}{(d+x)s+(d+1)x(1-s)}\leqslant\frac{(d+x)s}{(d+x)s+(d+x)(1-s)}=s.$
 By combining Corollary 4.2 with Lemma 4.4 we immediately get the following statement.
 Corollary 4.5.
-Let $H$ be a bipartite graph with minimum degree at least $\delta$. Then for $x\geq 2$ and $0\leq s\leq 1$ we have
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geq\prod_{v\in V}G(d_{v},x,s,s).$
+Let $H$ be a bipartite graph with minimum degree at least $\delta$. Then for $x\geqslant 2$ and $0\leqslant s\leqslant 1$ we have
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,s).$
 Proof.
 We have
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)\right]$
-$\geq\prod_{v\in V}\left[\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(\delta)\right)\right]$
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(\delta)\}_{u\in N(v)}\right)$
 $=\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))$
-$\geq\prod_{v\in V}G(d_{v},x,s,s).$
+$\geqslant\prod_{v\in V}G(d_{v},x,s,s).$
 Lemma 4.6.
-Let $x\geq 2$ and $\gamma\leq s<1$. A sufficient condition for
-$G(d,x,s,\gamma)\geq G(d+1,x,s,\gamma)$
-is $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq\gamma.$
-Furthermore, the function $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}$ is monotone increasing in $d$ for $x\geq 1$ and $d\geq 2$, so it is enough to check the above inequality for some fixed $d_{0}$ to conclude that the sequence $G(d,x,s,\gamma)$ is decreasing for $d\geq d_{0}$.
+Let $x\geqslant 1$ and $0\leqslant\gamma\leqslant s<1$. If $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geqslant\gamma$, then
+$G(d,x,s,\gamma)\geqslant G(d+1,x,s,0),$
+consequently
+$G(d,x,s,\gamma)\geqslant G(d+1,x,s,0)\geqslant G(d+1,x,s,\gamma).$
+Furthermore, the function $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}$ is monotone increasing in $d$ for $x\geqslant 1$ and $d\geqslant 2$, so it is enough to check the above inequality for some fixed $d_{0}$ to conclude that the sequence $G(d,x,s,\gamma)$ is decreasing for $d\geqslant d_{0}$.
 Proof.
-We have
-$G(d,x,s,\gamma)-G(d+1,x,s,\gamma)$
-$=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right)-\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)\left(s-\frac{s}{d+2}\gamma^{d+1}\right)$
-$=s^{2}\left(\frac{d+x}{d+1}-\frac{d+1+x}{d+2}\right)-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)+\frac{s}{d+2}\gamma^{d+1}\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)$
-$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
-$\geq\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}\gamma^{d}$
-which is non-negative if $\frac{x-1}{x(d+2)}\geq\gamma^{d-1}$. Finally, the derivative of $h(t):=\frac{1}{t-1}\ln\left(\frac{x-1}{x(t+2)}\right)$ is
+The inequality $G(d+1,x,s,0)\geqslant G(d+1,x,s,\gamma)$ is trivial. We only need to prove the inequality $G(d,x,s,\gamma)\geqslant G(d+1,x,s,0)$. We have
+$G(d,x,s,\gamma)-G(d+1,x,s,0)$
+$=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma^{d}\right)-\left(\frac{(d+1+x)s}{d+2}+x(1-s)\right)s$
+$=s^{2}\left(\frac{d+x}{d+1}-\frac{d+1+x}{d+2}\right)-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
+$=\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{s}{d+1}\gamma^{d}\left(\frac{(d+x)s}{d+1}+x(1-s)\right)$
+$\geqslant\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}\gamma^{d}$
+$\geqslant\frac{(x-1)s^{2}}{(d+1)(d+2)}-\frac{sx}{d+1}s\gamma^{d-1}$
+$=\frac{xs^{2}}{d+1}\left(\frac{x-1}{x(d+2)}-\gamma^{d-1}\right)$
+which is non-negative if $\frac{x-1}{x(d+2)}\geqslant\gamma^{d-1}$. Finally, the derivative of $h(t):=\frac{1}{t-1}\ln\left(\frac{x-1}{x(t+2)}\right)$ is
 $h^{\prime}(t)=\frac{1}{(t-1)^{2}}\ln\left(\frac{x(t+2)}{x-1}\right)-\frac{1}{(t-1)(t+2)}.$
 Then
-$h^{\prime}(t)\geq\frac{\ln(t+2)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}\geq\frac{\ln(3)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}>0$
+$h^{\prime}(t)\geqslant\frac{\ln(t+2)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}\geqslant\frac{\ln(3)}{(t-1)^{2}}-\frac{1}{(t-1)(t+2)}>0$
 if $t>1$. ∎
 5. Proof of Theorem 1.3.
-In this section we prove that
-$T_{M}(x,0)T_{M}(0,x)\geq T_{M}(1,1)^{2}$
-whenever $x\geq 2.355$. By the transfer lemma it is enough to prove that
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
-for any bipartite graph $H$ with minimum degree at least $1$. Note that $\widetilde{T}_{H}(1,1)=1$.
+In this section, we prove that
+$T_{M}(x,0)T_{M}(0,x)\geqslant T_{M}(1,1)^{2}$
+whenever $x\geqslant 2.355$. By the transfer lemma it is enough to prove that
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
+for any bipartite graph $H$ with minimum degree at least $1$. Recall that $\widetilde{T}_{H}(1,1)=1$.
 We do not immediately prove the best value $2.355$, but give various improvements to the value $2.9243$ step by step to digest one idea at a time.
@@
 Idea 1. By Corollary 4.5 we have
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(1)).$
-Since $\widetilde{T}_{H}(1,1)^{2}=1$ it is enough if $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for all $d\geq 1$. Let
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(1)).$
+Since $\widetilde{T}_{H}(1,1)^{2}=1$ it is enough if $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for all $d\geqslant 1$. Let
 $x=\frac{3+\sqrt{5}}{2}\approx 2.6180..\ \ \text{and}\ \ \ s=\frac{6+2\sqrt{5}}{7+3\sqrt{5}}\approx 0.7639$...
 These numbers are chosen in such a way that $G(1,x,s,\gamma_{x,s}(1))=1$.
-It is easy to check that for $d=9$ we have $\left(\frac{x-1}{x(d+2)}\right)^{1/(d-1)}\geq s\geq\gamma_{x,s}(1)$.
-By Lemma 4.6 this implies that for $d\geq 9$ we have $G(d,x,s,\gamma_{x,s}(1))\geq G(\infty,x,s)>1$.
-It is also easy to check (see the jupyter notebook or Table 1 in the Appendix) that $G(d,x,s,\gamma_{x,s}(1))\geq 1$ for $d\in\{2,\dots,10\}$.
+It is easy to check that for $d_{0}=9$ we have $\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}\geqslant s\geqslant\gamma_{x,s}(1)$.
+By Lemma 4.4 this implies that for $d\geqslant d_{0}$ we have $G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)>1$.
+It is also easy to check (see the jupyter notebook or Table 1 in the Appendix) that $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for $d\in\{2,\dots,8\}$.
 So with $x=\frac{3+\sqrt{5}}{2}$ we have
-$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geq\widetilde{T}_{H}(1,1)^{2}$
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)\geqslant\widetilde{T}_{H}(1,1)^{2}$
 implying that
-$T_{M}(x,0)T_{M}(0,x)\geq T_{M}(1,1)^{2}$
+$T_{M}(x,0)T_{M}(0,x)\geqslant T_{M}(1,1)^{2}$
 for all matroids without loops and coloops, in particular for all graphs $G$ without loops and bridges.
@@
 Observe that we can assume that $H$ is connected since $\widetilde{T}_{H}(x,y)=\prod_{i=1}^{k}\widetilde{T}_{H_{i}}(x,y)$ if $H$ has connected components $H_{1},\dots,H_{k}$.
-The statement $\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq 1$ is trivially true for $H=K_{2}$, the complete graph on $2$ vertices.
+The statement $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant 1$ is trivially true for $H=K_{2}$, the complete graph on $2$ vertices.
 So we can assume that $H$ has at least $3$ vertices.
 On the other hand, if $H$ is connected and has at least $3$ vertices, then a degree one vertex cannot be adjacent to another degree one vertex.
-Thus it is enough if
-$G(1,x,s,\gamma_{x,s}(2))\geq 1\ \ \text{and}\ \ G(d,x,s,\gamma_{x,s}(1))\geq 1$
-for all $d\geq 2$. This is satisfied if $x=2.54$ and $s=0.76$. (See Table 2.)
-Idea 3.
-Observe that it is not really necessary that $G(1,x,s,\gamma)\geq 1$, it is quite enough if for a non-leaf vertex the product with its leaf neighbors is at least $1$.
-If $x\geq 2.2$, then any vertex can have at most $2$ leaf neighbors, since for the star $S_{4}$ on $4$ vertices we have
-$\frac{1}{x}\widetilde{T}_{S_{4}}(x,0)\widetilde{T}_{S_{4}}(0,x)=\frac{1}{x}\cdot\frac{x^{3}+x^{2}+x}{4}\cdot\frac{x}{4}=\frac{x^{3}+x^{2}+x}{16}>1$
-so by the gluing lemma (Lemma 2.14) we know that a graph containing a pending $S_{4}$ cannot be a minimal counter example. So it is enough to check if for every $d\geq 2$ we have
-$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$
+We claim that it is enough to check that if
+(5.1) $G(1,x,s,\gamma_{x,s}(2))\geqslant 1\ \ \text{and}\ \ G(d,x,s,\gamma_{x,s}(1))\geqslant 1$
+for all $d\geqslant 2$. Indeed,
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$=\prod_{v\in V:d_{v}=1}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{v\in V:d_{v}\geqslant 2}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$\geqslant\prod_{v\in V:d_{v}=1}G(1,x,s,\gamma_{x,s}(2))\cdot\prod_{v\in V:d_{v}\geqslant 2}G(d_{v},x,s,\gamma_{x,s}(1))$
+$\geqslant 1.$
+We claim that the conditions of 5.1 are satisfied if $x=2.54$ and $s=0.76$, see Table 2.
+Clearly, the condition $G(1,x,s,\gamma_{x,s}(2))\geqslant 1$ is just a matter of computation.
+To prove that $G(d,x,s,\gamma_{x,s}(1))\geqslant 1$ for all $d\geqslant d_{0}=9$, we use the same argument as before: for $d\geqslant d_{0}$ the sequence is $G(d,x,s,\gamma_{x,s}(1))$ is monotone decreasing by Lemma 4.6 and checking that
+$\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}>\gamma_{x,s}(1).$
+This gives that
+$G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)>1$
+for all $d\geqslant d_{0}$. Since $G(d,x,s,\gamma_{x,s}(1))>1$ for $2\leqslant d\leqslant 8$ we get this inequality for all $d\geqslant 2$.
+Idea 3. The next idea is to group the degree $1$ vertices together with their neighbors:
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}\mathbb{G}\left(d_{u},x,s,\{\gamma_{x,s}(d_{u^{\prime}})\}_{u^{\prime}\in N(u)}\right)\right)$
+$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
+$\geqslant\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(1)\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
+$=\prod_{v\in V:d_{v}\geqslant 2}\left(G(d_{v},x,s,\gamma_{x,s}(1))\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
+It would be enough for us if for all $v$ with $d_{v}\geqslant 2$ we have
+$G(d_{v},x,s,\gamma_{x,s}(1))\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\geqslant 1.$
+In order to prove such an inequality, we need an upper bound for the number of degree $1$ neighbors of a vertex $v$.
+We claim that if $x\geqslant 2.2$, then any vertex can have at most $2$ leaf neighbors.
+To prove this, we use the gluing lemma.
+First, observe that for the star $S_{4}$ on $4$ vertices we have
+$\frac{1}{x}\widetilde{T}_{S_{4}}(x,0)\widetilde{T}_{S_{4}}(0,x)=\frac{1}{x}\cdot\frac{x^{3}+x^{2}+x}{4}\cdot\frac{x}{4}=\frac{x^{3}+x^{2}+x}{16}>1,$
+so by the gluing lemma (Lemma 2.13) we know that a graph containing a pendant $S_{4}$ cannot be a minimal counter example.
+This means that in a minimal counter example every vertex has at most $2$ neighbors of degree $1$.
+So, it is enough to check if for every $d\geqslant 2$ we have
+$\min_{0\leqslant k\leqslant\min(2,d-1)}G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{k}\geqslant 1.$
+Here we used that if $d=2$ then it cannot have two neighbors of degree $1$, since then $K_{1,2}$ would be a connected component of $H$, but it is clearly not a counter example.
+We show that for $x=2.36$ and $s=0.78$ these inequalities are satisfied.
+We can check this inequality separately for $d=2$.
+For $d>2$ we have
+$\ \min_{0\leqslant k\leqslant\min(2,d-1)}G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{k}$
+$=\min(G(d,x,s,\gamma_{x,s}(1)),G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{2}).$
+(Clearly, if $G(1,x,s,\gamma_{x,s}(d))>1$ then the minimum is achieved at the first term, otherwise the minimum is achieved at the second term.)
+In order to check that the above minimum is at least $1$ we proceed as before. Suppose that $G(\infty,x,s)\geqslant 1,$ and for some $d_{0}$ we have
+(5.2) $G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geqslant 1,$
+and for $d\geqslant d_{0}$ the sequence $G(d,x,s,\gamma_{x,s}(1))$ is monotone decreasing. To check this latter property, we only need to check
+$\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}>\gamma_{x,s}(1).$
+Then for $d\geqslant d_{0}$ we have
+$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{2}\geqslant G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geqslant 1,$
 and
-$G(d,x,s,\gamma_{x,s}(1))\geq 1.$
-In order to check the first inequality for large $d$ it is enough if for some $d_{0}$ we have
-$G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1,$
-then we get that
-$G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d))^{\min(2,d-1)}\geq 1$
-for $d\geq d_{0}$.
-So we only need to check the first inequality for $d=2,\dots,d_{0}$ and the inequality $G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{2}\geq 1$.
-These inequalities are satisfied for $x=2.36$, $s=0.78$ and $d_{0}=44$.
+$G(d,x,s,\gamma_{x,s}(1))\geqslant G(\infty,x,s)\geqslant 1.$
+With the choice $x=2.36,s=0.78$ and $d_{0}=44$ all these inequalities are satisfied.
+Then we can check the remaining inequalities for $d=3,\dots,43$ with a computer separately.
 Again see the jupyter notebook [7] or alternatively, Table 3.
-Idea 4. We can gain a very small further improvement as follows. For $d\geq 2$ we can use that
-$\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\prod_{u\in N(v)}\gamma_{x,s}(d_{u})\right)$
-$\geq\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\gamma_{x,s}(1)^{\min(2,d_{v}-1)}\gamma_{x,s}(2)^{d_{v}-\min(2,d_{v}-1)}\right),$
+Idea 4. We can gain a very small further improvement as follows. We start our series of inequalities just as in the previous step:
+$\widetilde{T}_{H}(x,0)\widetilde{T}_{H}(0,x)$ $\geqslant\prod_{v\in V}\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}\mathbb{G}\left(d_{u},x,s,\{\gamma_{x,s}(d_{u^{\prime}})\}_{u^{\prime}\in N(u)}\right)\right)$
+$=\prod_{v\in V:d_{v}\geqslant 2}\left(\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)\cdot\prod_{u\in N(v):d_{u}=1}G\left(1,x,s,\gamma_{x,s}(d_{v})\right)\right)$
+Previously, we bounded $\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$ with $G(d_{v},x,s,\gamma_{x,s}(1))$, but now we observe that if $d_{v}>2$, then $v$ can have at most two neighbors of degree $1$, while if $d_{v}=2$ then $v$ can have at most one vertex of degree $1$, all other neighbors have degree at least $2$.
+Thus, for $d_{v}\geq 2$ we can use that
+$\mathbb{G}\left(d_{v},x,s,\{\gamma_{x,s}(d_{u})\}_{u\in N(v)}\right)$
+$\geqslant\left(\frac{(d_{v}+x)s}{d_{v}+1}+x(1-s)\right)\left(s-\frac{s}{d_{v}+1}\gamma_{x,s}(1)^{\min(2,d_{v}-1)}\gamma_{x,s}(2)^{d_{v}-\min(2,d_{v}-1)}\right),$
 Let
-$G_{2}(d,x,s,\gamma_{1},\gamma_{2}):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma_{1}^{\min(2,d-1)}\gamma_{2}^{d-\min(2,d-1)}\right).$
-Then it is enough to check that
-$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\geq 1$
-for $d\geq 2$ and
-$G_{2}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))G(d,x,s,\gamma_{x,s}(1))^{\min(2,d-1)}\geq 1$
-for all $d\geq 2$. This leads to the tiny improvement $x=2.355$ ($s=0.78$ and $d=100$). See [7] or Table 4 in the Appendix for further details.
+$G_{k}(d,x,s,\gamma_{1},\gamma_{2}):=\left(\frac{(d+x)s}{d+1}+x(1-s)\right)\left(s-\frac{s}{d+1}\gamma_{1}^{k}\gamma_{2}^{d-k}\right).$
+Then it is enough to check that for all $d\geqslant 2$ we have
+$\min_{0\leqslant k\leqslant\min(2,d-1)}G_{k}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\cdot G(1,x,s,\gamma_{x,s}(d))^{k}\geqslant 1.$
+We show that this is indeed true if $x=2.355$ and $s=0.78$.
+First, we check it for $d\geqslant d_{0}=100$.
+By Lemma 4.6 the sequence $G(d,x,s,\gamma)$ is monotone decreasing for $d\geqslant d_{0}$ if $\left(\frac{x-1}{x(d_{0}+2)}\right)^{1/(d_{0}-1)}\geqslant\gamma$.
+Thus for $0\leqslant k\leqslant 2$ we have
+$G_{k}(d,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))\cdot G(1,x,s,\gamma_{x,s}(d))^{k}$ $\geqslant G(d,x,s,\gamma_{x,s}(1))G(1,x,s,\gamma_{x,s}(d_{0}))^{k}$
+$\geqslant G(\infty,x,s)G(1,x,s,\gamma_{x,s}(d_{0}))^{k}$
+$>1,$
+where the last inequality can be checked directly. For $d<d_{0}$ we again check the values by a computer. See [7] or Table 4 in the Appendix for further details.
+Remark 5.1.
+It turns out that the value $2.355$ can be further improved to $2.35$ by observing that one can disregard one particular inequality:
+$G_{2}(3,x,s,\gamma_{x,s}(1),\gamma_{x,s}(2))G(1,x,s,\gamma_{x,s}(3))^{2}\geqslant 1.$
+This inequality comes from the situation where a vertex $v$ of degree $3$ is adjacent to two vertices of degree $1$.
+But this means that $H$ can be obtained by gluing $H_{1}=S_{4}$ to some $H_{2}$ at one of its leaves.
+As we discussed at Idea 3 this would mean that $H$ is not a minimal counter example.
+Choosing $x=2.35$, $s=0.79$ and $d_{0}=1000$ the remaining parts of Idea 4 work verbatimely.
 6. Matroids with prescribed circuit lengths
@@
 Lemma 6.1.
-Let $k\geq 4$ be a real number. Let $H$ be a bipartite graph such that all degrees of $H$ lie in the interval $[k+1,k^{4}-2k^{2}-1]$. Then
-$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
+Let $k\geqslant 4$ be a real number. Let $H$ be a bipartite graph such that all degrees of $H$ lie in the interval $k+1,k^{4}-2k^{2}-1]$. Then
+$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\widetilde{T}_{H}(1,1)^{2}.$
 Proof.
 We will use that $\widetilde{T}_{H}(1,1)=1$, and Corollary 4.5:
-$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\prod_{v\in V}G(d_{v},2,s,s).$
+$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},2,s,s).$
 Note that
 $G(d,2,s,s)=\left(\frac{d+2}{d+1}\cdot s+2(1-s)\right)\left(s-\frac{s^{d+1}}{d+1}\right)=\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)\left(s-\frac{s^{d+1}}{d+1}\right).$
-We will show that if $s=1-\frac{1}{k^{2}}$, then for all $d\in[k+1,k^{4}-2k^{2}-1]$ we have $G(d,2,s,s)\geq 1$.
-First observe that $G(d,2,s,s)$ is a concave function of $s$ on the interval $(0,1)$ as its second derivative is
-$-\frac{(2-s)(d^{2}+d)s^{d}+2ds}{(d+1)s},$
-which is clearly negative.
-This means that $I_{d}:=\{s\ |\ G(d,2,s,s)\geq 1\}$ is an interval for each $d$.
+We will show that if $s=1-\frac{1}{k^{2}}$, then for all $d\in[k+1,k^{4}-2k^{2}-1]$ we have $G(d,2,s,s)\geqslant 1$.
+First, observe that $G(d,2,s,s)$ is a concave function of $s$ on the interval $(0,1)$ as its second derivative is
+$2\frac{d}{d+1}(1-s^{d})+\left(\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)\right)(-ds^{d-1})$
+which is clearly negative as both terms are negative.
+(Here we used the derivation rule $(h_{1}\cdot h_{2})^{\prime\prime}=h_{1}^{\prime\prime}\cdot h_{2}+2h_{1}^{\prime}\cdot h_{2}^{\prime}+h_{1}\cdot h_{2}^{\prime\prime}$ for $h_{1}=\frac{d+2}{d+1}+\frac{d}{d+1}(1-s)$ and $h_{2}=s-\frac{s^{d+1}}{d+1}$ with $h_{1}^{\prime}=-\frac{d}{d+1}$, $h_{1}^{\prime\prime}=0$, $h_{2}^{\prime}=1-s^{d}$, $h_{2}^{\prime\prime}=-ds^{d-1}$.)
+This means that
+$I_{d}:=\{s\ |\ G(d,2,s,s)\geqslant 1\}$
+is an interval for each $d$.
 We will show that $1-\frac{1}{k^{2}}\in I_{k+1}\cap I_{k^{4}-2k^{2}-1}$.
-By first choosing $k=d-1$ we get that $1-\frac{1}{(d-1)^{2}}\in I_{d}$.
+By first choosing k=d1k=d-1 we get that $1-\frac{1}{(d-1)^{2}}\in I_{d}$.
 If we choose $k^{2}=1+\sqrt{d+2}$, then $k^{4}-2k^{2}-1=d$ and so $1-\frac{1}{1+\sqrt{d+2}}\in I_{d}$.
 Hence $\left[1-\frac{1}{1+\sqrt{d+2}},1-\frac{1}{(d-1)^{2}}\right]\subseteq I_{d}$.
-This means that if $k+1\leq d\leq k^{4}-2k^{2}-1$, then $1-\frac{1}{k^{2}}\in I_{d}$.
-First we show that $1-\frac{1}{k^{2}}\in I_{k+1}$, that is, $G(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}})\geq 1$. We will use that
-$(1-t)^{r}\leq 1-\binom{r}{1}t+\binom{r}{2}t^{2}$
-for $0\leq t\leq 1$ by Bonferroni’s inequality applied to independent events with probability $t$. Hence
+This means that if $k+1\leqslant d\leqslant k^{4}-2k^{2}-1$, then $1-\frac{1}{k^{2}}\in I_{d}$.
+First we show that $1-\frac{1}{k^{2}}\in I_{k+1}$, that is, $G(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}})\geqslant 1$. We will use that
+$(1-t)^{r}\leqslant 1-\binom{r}{1}t+\binom{r}{2}t^{2}$
+for $0\leqslant t\leqslant 1$. This can be either proved by Bonferroni’s inequality applied to independent events with probability $t$, or by induction: the base case $r=1$ being trivial we have
+$(1-t)^{r}$ $=(1-t)^{r-1}(1-t)$
+$\leqslant\left(1-\binom{r-1}{1}t+\binom{r-1}{2}t^{2}\right)(1-t)$
+$=1-\binom{r}{1}t+\binom{r}{2}t^{2}-\binom{r-1}{2}t^{2}$
+$\leqslant 1-\binom{r}{1}t+\binom{r}{2}t^{2}.$
+Hence
 $G\left(d,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $=\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{1}{k^{2}}\right)^{d+1}\right)$
-$\geq\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{d+1}{k^{2}}+\frac{(d+1)d}{2k^{2}}\right)\right)$
+$\geqslant\left(\frac{d+2}{d+1}+\frac{d}{d+1}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{d+1}\left(1-\frac{d+1}{k^{2}}+\frac{(d+1)d}{2k^{4}}\right)\right).$
 For $d=k+1$ we get that
-$G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $\geq\left(\frac{k+3}{k+2}+\frac{k+1}{k+2}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{k+2}\left(1-\frac{k+2}{k^{2}}+\frac{(k+2)(k+1)}{2k^{2}}\right)\right)$
-$=1+\frac{1}{2}\frac{3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2}{k^{8}+4k^{7}+4k^{6}}$
-For $k\geq 4$ we have $03k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2>$ 0 showing that $G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)>1$.
+$G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)$ $\geqslant\left(\frac{k+3}{k+2}+\frac{k+1}{k+2}\cdot\frac{1}{k^{2}}\right)\left(1-\frac{1}{k^{2}}-\frac{1}{k+2}\left(1-\frac{k+2}{k^{2}}+\frac{(k+2)(k+1)}{2k^{4}}\right)\right)$
+$=1+\frac{1}{2}\frac{3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2}{k^{8}+4k^{7}+4k^{6}}.$
+For $k\geqslant 4$ we have $3k^{5}-4k^{4}-12k^{3}-10k^{2}-5k-2>0$ showing that $G\left(k+1,2,1-\frac{1}{k^{2}},1-\frac{1}{k^{2}}\right)>1$.
 Next we show that $1-\frac{1}{k^{2}}\in I_{k^{4}-2k^{2}-1}$. For $s=1-\frac{1}{k^{2}}$ and $d=k^{4}-2k^{2}-1$ we have
@@
 By $1-t<e^{-t}$ we have
-$\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}<e^{-\frac{1}{k^{2}}(k^{4}-2k^{2})}=e^{-k^{2}-2}.$
-Let us introduce $y=k^{2}-2$. Then $k^{6}-k^{4}-k^{2}-1=y^{3}+5y^{2}+7y+1$, so we need that
-$\frac{y}{y^{3}+5y^{2}+7y+1}<e^{-y}$
+$\left(1-\frac{1}{k^{2}}\right)^{k^{4}-2k^{2}}<e^{-\frac{1}{k^{2}}(k^{4}-2k^{2})}=e^{-k^{2}+2}.$
+Let us introduce $y=k^{2}-2$. Then $k^{6}-k^{4}-k^{2}-1=y^{3}+5y^{2}+7y+1$, so it is enough to prove that
+$\frac{y}{y^{3}+5y^{2}+7y+1}>e^{-y}$
 or equivalently
 $y^{2}+5y+7+\frac{1}{y}<e^{y}.$
-Since $k\geq 3$ we have $y=k^{2}-2\geq 7$, but $e^{y}>y^{2}+5y+7+\frac{1}{y}$ already true for $y\geq 4$.
-Remark 6.2.
-There was no attempt to optimize Lemma 6.1, not even concerning the minimum degree. It turns out that if $s=0.9226$, then
-$G(d,2,s,\gamma_{2,s}(3))>1$
-for $3\leq d\leq 141$. So if all degrees of a bipartite graph $H$ lies in the interval $[3,141]$, then
-$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
-Similarly, for $s=0.9622$ we have
-$G(d,2,s,\gamma_{2,s}(4))>1$
-for $4\leq d\leq 646$. So if all degrees of a bipartite graph $H$ lies in the interval $[4,646]$, then
-$\widetilde{T}_{H}(2,0)\widetilde{T}(0,2)\geq\widetilde{T}(1,1)^{2}.$
-This shows that in Theorem 1.4 we could have written $\ell\geq 4$ and probably a faster growing function than $(\ell-2)^{4}$.
+Since $k\geqslant 3$ we have $y=k^{2}-2\geqslant 7$, but $e^{y}>y^{2}+5y+7+\frac{1}{y}$ is already true for $y\geqslant 4$.
 Now we are ready to finish the proof of Theorem 1.4.
@@
 Let us consider the local basis exchange graph $H[A]$ on $A\cup B$.
-Observe that $v\in B$ are adjacent with exactly those elements that are in his fundamental circuit.
-So $\ell\leq d_{v}+1\leq(\ell-2)^{4}$.
-Similarly, if $u\in A$, then it is adjacent with those elements of $B$ that are in its fundamental circuits with respect to $B$ in $M^{*}$.
-So $\ell\leq d_{u}+1\leq(\ell-2)^{4}$.
-By choosing $k=\ell-1$ we get that $k^{4}-2k^{2}-1=(\ell-1)^{4}-2(\ell-1)^{2}-1\geq(\ell-2)^{4}$ if $\ell\geq 3$. So by Lemma 1.4 we have
-$\widetilde{T}_{H[A]}(2,0)\widetilde{T}_{H[A]}(0,2)\geq\widetilde{T}_{H[A]}(1,1)^{2}.$
-Since this is true for all basis $A$ of the matroid $M$, we get by the Transfer lemma (Lemma 2.6 and Remark 2.7) that
-$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
+Observe that $v\in B$ are adjacent with exactly those elements that are in its fundamental circuit.
+So $\ell\leqslant d_{v}+1\leqslant(\ell-2)^{2}(\ell^{2}-4\ell+2)$.
+Similarly, if $u\in A$, then it is adjacent with those elements of $B$ that are in its fundamental circuit with respect to $B$ in $M^{*}$.
+So $\ell\leqslant d_{u}+1\leqslant(\ell-2)^{2}(\ell^{2}-4\ell+2)$.
+We treat the cases $\ell=4$ and $5$ separately. If $\ell=4$, then all degree $d_{v}$ satisfy that $3\leqslant d_{v}\leqslant 7$. Let us choose $s=0.9226$. Then
+$G(d,2,s,\gamma_{2,s}(3))>1$
+for $3\leqslant d\leqslant 141$. By Corollary 4.5 we have
+$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(3))>1=\widetilde{T}_{H}(1,1)^{2}.$
+Together with the transfer lemma this implies the claim. In case of $\ell=5$ all degree $d_{v}$ satisfy $4\leqslant d_{v}\leqslant 62$. With the choice $s=0.9622$ we have
+$G(d,2,s,\gamma_{2,s}(4))>1$
+for $4\leqslant d\leqslant 646$. Hence again by Corollary 4.5 we have
+$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(\delta))\geqslant\prod_{v\in V}G(d_{v},x,s,\gamma_{x,s}(4))>1=\widetilde{T}_{H}(1,1)^{2}$
+and the transfer lemma this implies the claim.
+Now let us assume that $\ell\geqslant 6$. By choosing $k=\ell-2$ we get that $k\geq 4$ and for all vertex $v$ we have $k+1=\ell-1\leq d_{v}$ and
+$k^{4}-2k^{2}-1=(\ell-2)^{4}-2(\ell-2)^{2}-1\geqslant(d_{v}+1)-1=d_{v}.$
+This means that all degree $d_{v}$ satisfy $k+1\leqslant d_{v}\leqslant k^{4}-2k^{2}-1$. So by Lemma 6.1 we have
+$\widetilde{T}_{H[A]}(2,0)\widetilde{T}_{H[A]}(0,2)\geqslant\widetilde{T}_{H[A]}(1,1)^{2}.$
+Since this is true for all basis $A$ of the matroid $M$, we get by the Transfer lemma (Lemma 2.6) that
+$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
 7. Concluding remarks
-We end this paper with some conjectures.
+We end this paper with some conjectures. Let us start with an admittedly provocative conjecture.
+Conjecture 7.1.
+Let $x_{0}$ be the largest zero of the polynomial $x^{3}-9(x-1)$. Then for any matroid $M$ without loops and coloops we have
+$T_{M}(x_{0},0)T_{M}(0,x_{0})\geq T_{M}(1,1)^{2}.$
+We believe that even a disproof of this conjecture would lead to a better understanding of this problem than slightly improving the constant $2.355$ (or $2.35$).
+Concerning the Merino-Welsh conjecture, the following variant might be true.
+Conjecture 7.2.
+If $M$ is a simple co-simple matroid, then
+$T_{M}(2,0)T_{M}(0,2)\geqslant T_{M}(1,1)^{2}.$
+The following line of attack seems to be natural.
 Conjecture 7.3.
 If the bipartite graph $H$ has minimum degree at least $2$, then
-$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geq\widetilde{T}_{H}(1,1)^{2}.$
-If this conjecture is true, then the following conjecture would be an immediate corollary.
+$\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)\geqslant\widetilde{T}_{H}(1,1)^{2}.$
+Clearly, Conjecture 7.3 implies Conjecture 7.2.
+At this point, we must admit that we lack examples.
+Computational results suggest the following conjecture.
+(Here we only phrase the conjecture when the number of vertices is divisible by $4$, but there are similar conjectures when $n=4k+i$ for $i\in\{1,2,3\}$.)
 Conjecture 7.4.
-If $M$ is a simple co-simple matroid, then
-$T_{M}(2,0)T_{M}(0,2)\geq T_{M}(1,1)^{2}.$
-Acknowledgment. The author is very grateful to Csongor Beke, Gergely Kál Csáji and Sára Pituk for discussions on the topic of this paper.
+Let $n=4k$, then the following bipartite graph minimizes $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)$ among bipartite graphs on $n$ vertices without isolated vertices: add a leaf vertex to each vertex of a $K_{k,k}$.
+Figure 3. The conjectured minimizer graph for $\widetilde{T}_{H}(2,0)\widetilde{T}_{H}(0,2)$ on $24$ vertices.
+If this conjecture is true, then the Merino-Welsh conjecture is true for matroids on $n$ elements when $n\leqslant 40$ and divisible by $4$, which shows why it is hard to find a counter example for the Merino-Welsh conjecture just by brute force computer search.
+Acknowledgment.
+The author is very grateful to Csongor Beke, Ferenc Bencs, Gergely Kál Csáji and Sára Pituk for discussions on the topic of this paper.
+The author also thanks the very careful work of the two referees, they spotted many typos and gave very useful suggestions.